### v1

# Reinforced Diffuser for Red Teaming Large Vision-Language Models

Abstract
The rapid advancement of large Vision-Language Models (VLMs) has raised significant safety concerns, particularly regarding their vulnerability to jailbreak attacks. While existing research primarily focuses on VLMs’ susceptibility to harmful instructions, this work identifies a critical yet overlooked vulnerability: current alignment mechanisms often fail to address the risks posed by toxic text continuation tasks. To investigate this issue, we propose a novel Red Team Diffuser (RTD) framework, which leverages reinforcement learning to generate red team images that effectively induce highly toxic continuations from target black-box VLMs. The RTD pipeline begins with a greedy search for high-quality image prompts that maximize the toxicity of VLM-generated sentence continuations, guided by a Large Language Model (LLM). These prompts are then used as input for the reinforcement fine-tuning of a diffusion model, which employs toxicity and alignment rewards to further amplify harmful outputs. Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% on the original attack set and 8.91% on a hold-out set. Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA. These findings reveal significant deficiencies in existing alignment strategies, particularly their inability to prevent harmful continuations. Our work underscores the urgent need for more robust and adaptive alignment mechanisms to ensure the safe deployment of VLMs in real-world applications.
Disclaimer: This paper contains content that may be disturbing or offensive.
footnotetext: $\dagger$  Corresponding author.
1 Introduction
The widespread deployment of large Vision-Language Models (VLMs), such as OpenAI’s GPT-4o Achiam et al. (2023) and Google’s Gemini Team et al. (2023), has raised significant concerns about their safety to potential jailbreak attacks. While VLMs inherit alignment properties from Large Language Models (LLMs), their multimodal nature introduces novel safety risks. Current jailbreak attacks on VLMs leverage either unimodal or multimodal approaches to manipulate the models into responding to harmful requests, such as providing bomb-making instructions Wang et al. (2024a).
However, research remains limited on the toxicity of generated outputs, such as insults and offensive language. For responsible deployment, it is crucial to rigorously safeguard these models against producing any harmful or unsafe content.
In this work, we demonstrate that existing safety mechanisms fail to prevent the generation of offensive, discriminatory, or hateful language when VLMs encounter biased or misleading contexts in text continuation tasks. In such scenarios, they often produce continuations containing harmful expressions, exposing a critical vulnerability.
Figure 1: Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as Stable Diffusion.
As multimodal models, the generated content of VLMs can be significantly influenced by visual modalities, despite having the same text input. Identifying which images can effectively serve as triggers for toxic continuations remains an open challenge. Building on the substantial progress made in white-box jailbreaks Qi et al. (2024); Wang et al. (2024a), one potential approach for triggering toxic continuation is the use of adversarial images. However, optimization-based adversarial images require full access to VLM parameters, which is impractical in real-world applications, especially in VLM-as-a-Service scenarios where the VLM operates as a black-box system. Additionally, such adversarial images often exhibit noticeable distortions, making them susceptible to detection and defense mechanisms Nie et al. (2022); Zhang et al. (2023); Zhao et al. (2024). Moreover, studies have demonstrated that adversarial images typically show limited transferability between VLMs, even when crafted to target multiple VLMs simultaneously Schaeffer et al. .
These limitations motivate us to explore alternative black-box attack strategies that do not rely on full model access and generate less detectable perturbations. Current black-box methods typically rely on manually crafted pipelines to generate image-text pairs, often involving typographic attacks and the use of pre-trained diffusion models to generate query-relevant images Gong et al. (2023); Liu et al. (2023); Teng et al. (2024). However, these approaches suffer from several drawbacks. First, structure-based attacks result in a restricted attack surface, producing predictable patterns that can be easily mitigated by defense mechanisms Wang et al. (2024c). Second, such images are primarily designed to bypass sensitive word detection and are not well-suited for triggering harmful continuations in our task. Recently, automated red-teaming approaches for VLMs have gained attention. IDEATOR Wang et al. (2024b) introduced the concept of “VLMs against VLMs”, employing a red team VLM to generate jailbreak text prompts paired with images created using pre-trained diffusion models Rombach et al. (2022). While this represents significant progress, it raises critical questions about the effectiveness of pre-trained diffusion models in jailbreaking tasks. Are these pre-trained models truly optimal for inducing harmful content in VLMs? Can images generated directly by diffusers reliably trigger such behavior? To answer these questions, we explore the potential of reinforced diffusers, which leverage reinforcement fine-tuning to optimize red team image generation for more effective and targeted VLM jailbreaking.
In this paper, we propose Red Team Diffuser (RTD), a novel diffusion model framework designed to generate unrestricted images that effectively bypass safety mechanisms in VLMs. Our approach leverages a two-stage process to systematically exploit vulnerabilities in VLMs. First, we utilize an LLM to perform a greedy search for image prompts that maximize the toxicity of VLM responses, identifying highly effective image prompts for red teaming purposes. Second, we fine-tune a diffusion model through reinforcement learning, jointly optimizing a toxicity reward to amplify the harmfulness of generated outputs and an image-prompt alignment reward to ensure consistency between prompts and images as a regularization term. Figure 1 illustrates the effectiveness of our proposed method in attacking LLaVA Liu et al. (2024b), showcasing its ability to systematically induce harmful outputs. This framework reveals critical vulnerabilities in VLMs, underscoring the need for further research to address these safety challenges.
The main contributions of our work can be summarized as:
  • We introduce Red Team Diffuser (RTD), a novel framework that, to the best of our knowledge, is the first to investigate the use of unrestricted images to effectively induce harmful outputs from VLMs.
  • We propose a novel two-stage black-box attack strategy, First, an LLM performs a greedy search to identify image prompts that maximize the toxicity of VLM responses. Second, a diffusion model is fine-tuned through reinforcement learning, leveraging toxicity and alignment rewards to enhance harmful outputs.
  • Experiments demonstrate that RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set. Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, as well as toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s Llama.
2 Related Work
2.1 Large Vision-Language Models
Large Vision-Language Models (VLMs) extend the functionality of Large Language Models (LLMs) by integrating visual modalities via an image encoder and modality alignment mechanisms. For example, MiniGPT-4 Zhu et al. (2023) aligned a frozen vision encoder Dosovitskiy et al. (2020) with a frozen Vicuna Chiang et al. (2023) using a single linear projection layer. InstructBLIP Dai et al. (2023), based on BLIP-2 Li et al. (2023b), introduced an instruction-tuning paradigm that leverages additional instructions to extract more useful visual features. LLaVA Liu et al. (2024b) used the text-only GPT-4 Achiam et al. (2023) to generate multimodal text-image instruction-following data and served as an end-to-end trained model that connects a vision encoder Radford et al. (2021) with LLaMA Touvron et al. (2023). While these models demonstrate significant potential in multimodal tasks, the integration of an additional visual modality introduces new avenues for attacks, distinct from those in text-only LLMs Shayegani et al. (2023b); Liu et al. (2024a); Li et al. (2023a); Ma et al. (2025).
Figure 2: Overview of the proposed method. Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation. The diffusion model generates an image based on this prompt. Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
2.2 Jailbreak Attacks against VLMs
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks Bagdasaryan et al. (2023); Bailey et al. (2023); Shayegani et al. (2023a); Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024); Wang et al. (2024a) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, making them less practical and more easily detectable Nie et al. (2022); Zhang et al. (2023). In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR. For example, FigStep Gong et al. (2023) converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP Liu et al. (2023); Ma et al. (2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems. To address these limitations, the concept of red team VLMs has advanced with IDEATOR Wang et al. (2024b), which autonomously generates jailbreak image-text pairs in a black-box manner. By providing a scalable and efficient solution for creating adversarial inputs, IDEATOR has demonstrated the potential of red teaming to expose vulnerabilities in VLMs. However, while IDEATOR focuses on leveraging red team VLMs, our work broadens the scope of red teaming by exploring the feasibility of red team diffusers, which are orthogonal to red team VLMs and pose greater challenges. Specifically, we investigate the capability of unrestricted images to trigger harmful content generation, highlighting the significant threats posed by such inputs to VLMs. Compared to existing approaches, our work presents a more challenging scenario by addressing the limitations of typography-based attacks and provides a rigorous framework for evaluating the robustness of VLMs against adversarial inputs.
3 Preliminary
Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm Black et al. (2023) to fine-tune a pretrained diffusion model Rombach et al. (2022). Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning (RL) to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP). At each step, actions (denoising predictions) are chosen to maximize the cumulative reward. The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps. Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$. The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process. The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$. The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$. That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
The goal of DDPO is to maximize the expected cumulative reward, which is defined over the sample distribution $p_{\theta}(x_{0}\mid c)$ and the context distribution $p(c)$:
$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right].$
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator Kakade and Langford (2002):
$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg{[}\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg{]},$
where the expectation is taken over denoising trajectories generated by the previous policy $\pi_{\theta_{\text{old}}}$. This estimator leverages the ratio of the current model’s likelihood to the previous model’s likelihood, allowing for more efficient updates.
4 Methodology
4.1 Threat Model
Attack Goals  We focus on the sentence continuation task, an underexplored area in alignment mechanisms that poses unique challenges due to the open-ended nature of model outputs. The attacker aims to bypass the model’s safety measures and induce toxic continuations, such as the generation of obscene or profane language. Toxic continuations can also be viewed as a form of multimodal prompt injection, where carefully designed text prefixes and image triggers manipulate the model’s output into generating harmful content.
Adversary Capabilities  We assume the attacker has only black-box access to the victim VLM, simulating real-world interactions via external interfaces. By reinforcement fine-tuning red team diffusers based on the VLM’s feedback, RTD learns the VLM’s behavior patterns and identifies vulnerabilities, refining its policy to trigger toxic content.
4.2 Greedy Prompt Search
A central challenge in our task is the design of effective image prompts to trigger harmful content. To address this challenge and integrate the toxic continuation task into a multimodal framework, we utilize Google’s Gemini Team et al. (2023), disabling its safety settings to generate image prompts for each given sentence. Specifically, we instruct Gemini to produce image prompts that preserve the original sentence’s semantics while amplifying harmful themes, thereby increasing the likelihood of generating toxic content. Additionally, we leverage in-context learning Brown (2020) by providing examples of previously identified harmful sentence continuations to guide Gemini in generating image prompts consistent with these harmful patterns.
However, a key limitation arises from the fact that image prompts generated by LLMs may not be optimal for eliciting highly toxic continuations. To address this, we propose a greedy prompt search algorithm designed to identify near-optimal image prompts. For each sentence, the algorithm iteratively generates image prompts using Gemini and evaluates their associated toxicity scores. The toxicity score of each new prompt is compared with the previous one: if the new score is higher, the iteration continues; otherwise, the search terminates.
4.3 Reward Functions for RTD
We now introduce the dual-reward mechanism for RTD, which consists of a toxicity reward and an alignment reward.
Toxicity Reward: In the toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, which is paired with an image $\mathbf{x}$ generated by the RTD. The toxicity of the generated continuation is then assessed using the Detoxify classifier Hanu and Unitary team (2020), which evaluates toxicity across six attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). The average of these scores is used as the reward function to guide the diffusion model’s optimization process:
$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$
Alignment Reward: Following Black et al. (2023), we introduce an alignment reward based on BERTScore Zhang et al. (2019) to improve multimodal understanding. In the context of our black-box attack, we expect the generated image $\mathbf{x}$ to be correctly interpreted by the target VLM and aligned with the intended semantics of the given image prompt $p$. Furthermore, we treat this alignment reward as a regularization term, helping to mitigate the risk of model collapse. Specifically, we define the alignment reward as:
$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big{(}S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big{)}$
[…]
Total Reward: The total reward is a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$
This combined reward allows the RTD to generate both toxic and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
Algorithm 1 Training Procedure of Red Team Diffuser
Require: Pre-trained Diffuser $D_{\theta}$, frozen LLM (e.g., Gemini) $\mathcal{G}$, target VLM $\mathcal{M}$, sentence continuation corpus $Y:=\{\mathbf{y}^{i}\}_{i=1}^{m}$, batch size $b$, learning rate $\alpha$, policy update steps $N$;
1:  Initialize $D_{\theta}$ with pre-trained weights and $P=\{\}$.
2:  for each sentence $\mathbf{y}^{i}\in Y$ do
3:     Initialize toxicity score $\mathcal{T}^{\text{max}}=0$.
4:     Query $\mathcal{G}$ with $\mathbf{y}^{i}$ to generate initial image prompt $p^{i}$.
5:     Generate initial image $\mathbf{x}^{i}=D_{\theta}(p^{i})$.
6:     Query $\mathcal{M}$ with $(\mathbf{x}^{i},\mathbf{y}^{i})$ to compute toxicity score $\mathcal{T}^{i}$.
7:     while $\mathcal{T}^{i}>\mathcal{T}^{\text{max}}$ do
8:        Update $\mathcal{T}^{\text{max}}=\mathcal{T}^{i}$ and store current prompt $p^{i}$.
9:        Query $\mathcal{G}$ with $\mathbf{y}^{i}$ to generate a new prompt $p^{i}$.
10:        Generate new image $\mathbf{x}^{i}=D_{\theta}(p^{i})$.
11:        Query $\mathcal{M}$ with $(\mathbf{x}^{i},\mathbf{y}^{i})$ to compute updated toxicity score $\mathcal{T}^{i}$.
12:     end while
13:     Append the stored best prompt $p^{i}$ into $P$.
14:  end for
15:  for $step=1,...,N$ do
16:     Sample a batch of sentence-prompt pairs $\{(\mathbf{y}^{k},p^{k})\}_{k=1}^{b}$ from $Y$ and $P$.
17:     for each $(\mathbf{y}^{k},p^{k})$ in the batch do
18:        Generate image $\mathbf{x}^{k}=D_{\theta}(p^{k})$.
19:        Query $\mathcal{M}$ with $(\mathbf{x}^{k},\mathbf{y}^{k})$ to compute total reward $\mathcal{R}_{\text{total}}$.
20:     end for
21:     Update the Diffuser $D_{\theta}$ using the policy gradient: $\theta\leftarrow\theta+\alpha\cdot\nabla_{\theta}J$.
22:  end for
23:  return Fine-tuned Diffuser $D_{\theta}^{*}$.
5 Experiments
Table 1: Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
No Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
Base Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
Greedy attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
RTD w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
RTD w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
RTD w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
RTD w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
No Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
Base Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
Greedy attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
RTD w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
RTD w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
RTD w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
RTD w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
Datasets We use the RealToxicityPrompts benchmark Gehman et al. (2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations Schick et al. (2021); Mehrabi et al. (2022); Qi et al. (2024). We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
Metric: Toxicity Rate To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark Gehman et al. (2020), we utilize the Detoxify classifier Hanu and Unitary team (2020) during training, and both Detoxify and the Perspective API during testing. Both classifiers assess toxicity across six predefined attributes. Each attribute’s toxicity score ranges from 0 (least toxic) to 1 (most toxic). The Perspective API is employed during testing to assess cross-classifier transferability. For each attribute, we calculate the proportion of continuations whose toxicity scores exceed a 0.5 threshold.
Implementation Details We use Stable Diffusion v1.5 Rombach et al. (2022) as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation. During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU. We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
5.1 Main Results
To evaluate the effectiveness of the Red Team Diffuser (RTD), we conduct black-box attacks on the original attack set and cross-prompt transfer attacks on the hold-out set against LLaVA. We compare several attack strategies: Base Attack: Gemini generates a single image prompt without optimization, providing a baseline for comparison. Greedy Attack: We iteratively refine the image prompt using the proposed greedy prompt search to maximize toxicity, selecting the most effective prompt at each step. RTD: We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
Table 2: Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
No Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
Base Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
Greedy attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
RTD w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
RTD w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
RTD w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
RTD w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
No Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
Base Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
Greedy attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
RTD w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
RTD w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
RTD w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
RTD w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
Original Attack Set Table 1 presents the results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the outputs, with ‘Any*’ indicating whether the text exhibits at least one of the six toxic attributes. The proposed greedy prompt search demonstrates a significant increase in toxicity compared to the Base Attack. For example, using the Detoxify classifier, the Greedy Attack achieved an ‘Any*’ rate of 64.12%, compared to 55.72% for the Base Attack. This substantial improvement underscores the effectiveness of greedy prompt search in maximizing toxic content generation. Furthermore, the Red Team Diffuser (RTD) consistently outperforms the Greedy Attack, with RTD w/ alignment (M) achieving the highest ‘Any*’ rate of 68.70% according to Detoxify. This highlights the considerable impact of reinforced fine-tuning in enhancing the RTD’s ability to trigger toxic continuations. Although RTD was trained using Detoxify, the results from the Perspective API validate its cross-classifier transferability. For example, RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% on Perspective API, demonstrating the robustness of RTD across different toxicity evaluation models.
Table 3: Cross-model transfer attack results on Gemini-1.5-flash. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
(%) Detoxify Perspective API
No Attack 76.23 78.26
RTD w/o alignment 80.30 81.49
RTD w/ alignment (L) 79.14 80.83
RTD w/ alignment (M) 81.33 82.52
RTD w/ alignment (H) 81.00 82.11
Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Importantly, the prompts in the hold-out set were not seen during training, making this evaluation a strong indicator of the generalization capability of the proposed methods. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy Attack demonstrates a clear improvement over the Base Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Base Attack. Similarly, with the Perspective API, the Greedy Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Base Attack. The Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These results clearly validate the effectiveness of our RTD in eliciting toxic continuations from diverse, unseen prompts. The consistent performance across both classifiers demonstrates its strong generalization capability beyond the training data.
Table 4: Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
(%) Detoxify Perspective API
No Attack 15.02 18.68
RTD w/o alignment 41.01 43.90
RTD w/ alignment (L) 41.85 45.24
RTD w/ alignment (M) 41.81 44.37
RTD w/ alignment (H) 41.15 43.78
5.2 Cross-Model Transferability
We evaluate the cross-model transferability of the proposed Red Team Diffuser (RTD) by using the RTD trained on LLaVA Liu et al. (2024b). Transfer attacks are performed on two additional VLMs: Gemini-1.5-flash Team et al. (2023) and Llama-3.2-11B-Vision-Instruct. Tables 3 and 4 present the ‘Any*’ rates evaluated by both the Detoxify classifier and the Perspective API. These results demonstrate that RTD, trained specifically on a single VLM, can effectively trigger toxic content when applied to unseen models, underscoring its robustness and transferability across VLMs. Table 3 summarizes the attack performance on Gemini 1.5 Flash. RTD w/ alignment (M) achieves the highest ‘Any*’ rates of 81.33% according to Detoxify and 82.52% according to the Perspective API, significantly outperforming other configurations. This indicates that the RTD maintains strong transferability to Gemini-1.5-flash, even without further fine-tuning. Table 4 presents the results for Llama 3.2 11B Vision-Instruct, a VLM with more advanced alignment mechanisms. While the baseline (No Attack) exhibits relatively low toxicity rates (15.02% according to Detoxify and 18.68% according to Perspective API), the RTD significantly increases these metrics. Notably, the lower baseline toxicity of Llama 3.2 can be attributed to its high refusal rate for toxic continuation tasks, demonstrating a stronger alignment mechanism that often declines requests to generate potentially harmful content. In contrast, other VLMs, such as LLaVA and Gemini, rarely refuse these requests, leading to higher baseline toxicity levels. Despite this, RTD is able to effectively bypass Llama 3.2’s refusal mechanisms, significantly increasing toxicity rates. Specifically, RTD w/ alignment (L) achieves the highest ‘Any*’ rates of 41.85% according to Detoxify and 45.24% according to Perspective API, closely followed by RTD w/ alignment (M). These findings demonstrate that RTD effectively transfers its attack capabilities to VLMs with different architectures and scales, even when those models have stronger alignment mechanisms.
5.3 Visualizations
Figure 3 shows a qualitative comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (RTD) (right). For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2. Despite being given the same image prompts, the outputs from the two models differ significantly. While the pre-trained Stable Diffusion produces neutral and benign visuals, RTD generates images that subtly evoke discomfort or unease. For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features illustrate how RTD effectively adapts image features to exploit vulnerabilities in VLMs, thereby enhancing its ability to induce harmful continuations.
Figure 3: Comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (right). Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
6 Limitations
The RTD framework demonstrates strong performance in eliciting harmful outputs from VLMs, but several limitations deserve attention. First, this study focuses exclusively on toxic continuation tasks. Broader challenges, such as misinformation propagation and ethical biases in complex multimodal scenarios, remain unexplored. Second, RTD’s high computational demands, including the use of 8 NVIDIA A100 GPUs, may hinder its adoption by researchers with limited resources. Finally, RTD poses ethical risks, as its methods could potentially be misused by adversaries.
7 Conclusion
In this work, we propose the Red Team Diffuser (RTD) framework to identify and exploit vulnerabilities in VLMs, specifically in toxic continuation tasks. Leveraging greedy prompt search and reinforcement fine-tuning, RTD generates images that effectively trigger toxic continuations, exposing the limitations of existing alignment mechanisms. Experimental results validate the effectiveness of RTD, showing significant increases in toxicity rates across both black-box and cross-model settings. For instance, RTD raises LLaVA’s toxicity rates by 10.69% on the original attack set and 8.91% on a hold-out set, while also demonstrating strong cross-model transferability, with increases of 5.1% on Gemini and 26.83% on Llama-3.2-11B-Vision-Instruct. Qualitative analysis further highlights RTD’s ability to subtly manipulate visual features, generating images that are more provocative and capable of inducing harmful continuations. These findings highlight critical gaps in current alignment strategies and emphasize the need for more robust and comprehensive mechanisms to enhance the safety of VLMs in real-world applications.

### v2

@@
-Reinforced Diffuser for Red Teaming Large Vision-Language Models
-Abstract
-The rapid advancement of large Vision-Language Models (VLMs) has raised significant safety concerns, particularly regarding their vulnerability to jailbreak attacks.
-While existing research primarily focuses on VLMs’ susceptibility to harmful instructions, this work identifies a critical yet overlooked vulnerability: current alignment mechanisms often fail to address the risks posed by toxic text continuation tasks.
-To investigate this issue, we propose a novel Red Team Diffuser (RTD) framework, which leverages reinforcement learning to generate red team images that effectively induce highly toxic continuations from target black-box VLMs.
-The RTD pipeline begins with a greedy search for high-quality image prompts that maximize the toxicity of VLM-generated sentence continuations, guided by a Large Language Model (LLM).
-These prompts are then used as input for the reinforcement fine-tuning of a diffusion model, which employs toxicity and alignment rewards to further amplify harmful outputs.
-Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% on the original attack set and 8.91% on a hold-out set.
+Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning
+Abstract.
+The growing deployment of large Vision-Language Models (VLMs) exposes critical safety gaps in their alignment mechanisms.
+While existing jailbreak studies primarily focus on VLMs’ susceptibility to harmful instructions, we reveal a fundamental yet overlooked vulnerability: toxic text continuation – where VLMs produce highly toxic completions when prompted with harmful text prefixes paired with semantically adversarial images.
+To systematically study this threat, we propose Red Team Diffuser (RTD), the first red teaming diffusion model that coordinates adversarial image generation and toxic continuation through reinforcement learning.
+Our key innovations include dynamic cross-modal attack and stealth-aware optimization.
+For toxic text prefixes from an LLM safety benchmark, we conduct greedy search to identify optimal image prompts that maximally induce toxic completions.
+The discovered image prompts then drive RL-based diffusion model fine-tuning, producing semantically-aligned adversarial images that boost toxicity rates.
+Stealth-aware optimization introduces joint adversarial rewards that balance toxicity maximization (via Detoxify classifier) and stealthiness (via BERTScore), circumventing traditional noise-based adversarial patterns.
+Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% over text-only baselines on the original attack set and 8.91% on unseen set, proving generalization capability.
 Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA.
-These findings reveal significant deficiencies in existing alignment strategies, particularly their inability to prevent harmful continuations.
-Our work underscores the urgent need for more robust and adaptive alignment mechanisms to ensure the safe deployment of VLMs in real-world applications.
+Our findings expose two critical flaws in current VLM alignment: (1) failure to prevent toxic continuation from harmful prefixes, and (2) overlooking cross-modal attack vectors.
+These results necessitate a paradigm shift toward multimodal red teaming in safety evaluations.
+We will publicly release the RTD framework to facilitate future research on robust VLM alignment.
 Disclaimer: This paper contains content that may be disturbing or offensive.
-footnotetext: $\dagger$  Corresponding author.
-1 Introduction
-The widespread deployment of large Vision-Language Models (VLMs), such as OpenAI’s GPT-4o Achiam et al.
-(2023) and Google’s Gemini Team et al.
-(2023), has raised significant concerns about their safety to potential jailbreak attacks.
-While VLMs inherit alignment properties from Large Language Models (LLMs), their multimodal nature introduces novel safety risks.
-Current jailbreak attacks on VLMs leverage either unimodal or multimodal approaches to manipulate the models into responding to harmful requests, such as providing bomb-making instructions Wang et al.
-(2024a).
-However, research remains limited on the toxicity of generated outputs, such as insults and offensive language.
-For responsible deployment, it is crucial to rigorously safeguard these models against producing any harmful or unsafe content.
-In this work, we demonstrate that existing safety mechanisms fail to prevent the generation of offensive, discriminatory, or hateful language when VLMs encounter biased or misleading contexts in text continuation tasks.
-In such scenarios, they often produce continuations containing harmful expressions, exposing a critical vulnerability.
-Figure 1: Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as Stable Diffusion.
-As multimodal models, the generated content of VLMs can be significantly influenced by visual modalities, despite having the same text input.
-Identifying which images can effectively serve as triggers for toxic continuations remains an open challenge.
-Building on the substantial progress made in white-box jailbreaks Qi et al.
-(2024); Wang et al.
-(2024a), one potential approach for triggering toxic continuation is the use of adversarial images.
-However, optimization-based adversarial images require full access to VLM parameters, which is impractical in real-world applications, especially in VLM-as-a-Service scenarios where the VLM operates as a black-box system.
-Additionally, such adversarial images often exhibit noticeable distortions, making them susceptible to detection and defense mechanisms Nie et al.
-(2022); Zhang et al.
-(2023); Zhao et al.
-(2024).
-Moreover, studies have demonstrated that adversarial images typically show limited transferability between VLMs, even when crafted to target multiple VLMs simultaneously Schaeffer et al.
-.
-These limitations motivate us to explore alternative black-box attack strategies that do not rely on full model access and generate less detectable perturbations.
-Current black-box methods typically rely on manually crafted pipelines to generate image-text pairs, often involving typographic attacks and the use of pre-trained diffusion models to generate query-relevant images Gong et al.
-(2023); Liu et al.
-(2023); Teng et al.
-(2024).
-However, these approaches suffer from several drawbacks.
-First, structure-based attacks result in a restricted attack surface, producing predictable patterns that can be easily mitigated by defense mechanisms Wang et al.
-(2024c).
-Second, such images are primarily designed to bypass sensitive word detection and are not well-suited for triggering harmful continuations in our task.
-Recently, automated red-teaming approaches for VLMs have gained attention.
-IDEATOR Wang et al.
-(2024b) introduced the concept of “VLMs against VLMs”, employing a red team VLM to generate jailbreak text prompts paired with images created using pre-trained diffusion models Rombach et al.
-(2022).
-While this represents significant progress, it raises critical questions about the effectiveness of pre-trained diffusion models in jailbreaking tasks.
-Are these pre-trained models truly optimal for inducing harmful content in VLMs?
-Can images generated directly by diffusers reliably trigger such behavior?
-To answer these questions, we explore the potential of reinforced diffusers, which leverage reinforcement fine-tuning to optimize red team image generation for more effective and targeted VLM jailbreaking.
-In this paper, we propose Red Team Diffuser (RTD), a novel diffusion model framework designed to generate unrestricted images that effectively bypass safety mechanisms in VLMs.
-Our approach leverages a two-stage process to systematically exploit vulnerabilities in VLMs.
-First, we utilize an LLM to perform a greedy search for image prompts that maximize the toxicity of VLM responses, identifying highly effective image prompts for red teaming purposes.
-Second, we fine-tune a diffusion model through reinforcement learning, jointly optimizing a toxicity reward to amplify the harmfulness of generated outputs and an image-prompt alignment reward to ensure consistency between prompts and images as a regularization term.
-Figure 1 illustrates the effectiveness of our proposed method in attacking LLaVA Liu et al.
-(2024b), showcasing its ability to systematically induce harmful outputs.
-This framework reveals critical vulnerabilities in VLMs, underscoring the need for further research to address these safety challenges.
-The main contributions of our work can be summarized as:
-  • We introduce Red Team Diffuser (RTD), a novel framework that, to the best of our knowledge, is the first to investigate the use of unrestricted images to effectively induce harmful outputs from VLMs.
-  • We propose a novel two-stage black-box attack strategy, First, an LLM performs a greedy search to identify image prompts that maximize the toxicity of VLM responses.
-Second, a diffusion model is fine-tuned through reinforcement learning, leveraging toxicity and alignment rewards to enhance harmful outputs.
-  • Experiments demonstrate that RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set.
-Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, as well as toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s Llama.
-2 Related Work
-2.1 Large Vision-Language Models
-Large Vision-Language Models (VLMs) extend the functionality of Large Language Models (LLMs) by integrating visual modalities via an image encoder and modality alignment mechanisms.
-For example, MiniGPT-4 Zhu et al.
-(2023) aligned a frozen vision encoder Dosovitskiy et al.
-(2020) with a frozen Vicuna Chiang et al.
-(2023) using a single linear projection layer.
-InstructBLIP Dai et al.
-(2023), based on BLIP-2 Li et al.
-(2023b), introduced an instruction-tuning paradigm that leverages additional instructions to extract more useful visual features.
-LLaVA Liu et al.
-(2024b) used the text-only GPT-4 Achiam et al.
-(2023) to generate multimodal text-image instruction-following data and served as an end-to-end trained model that connects a vision encoder Radford et al.
-(2021) with LLaMA Touvron et al.
-(2023).
-While these models demonstrate significant potential in multimodal tasks, the integration of an additional visual modality introduces new avenues for attacks, distinct from those in text-only LLMs Shayegani et al.
-(2023b); Liu et al.
-(2024a); Li et al.
-(2023a); Ma et al.
-(2025).
-Figure 2: Overview of the proposed method.
+Red Teaming, Vision-Language Models, Diffusion Models, Reinforcement Learning
+ccs: Computing methodologies Artificial intelligence
+1. Introduction
+Figure 1.
+Example of a successful jailbreak on LLaVA.
+The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as general-purpose diffusion model (Rombach et al., 2022).
+The rapid deployment of large Vision-Language Models (VLMs), such as GPT-4o (Achiam et al., 2023) and Gemini (Team et al., 2023), has exposed critical blind spots in their multimodal alignment mechanisms.
+While existing safety strategies predominantly focus on preventing responses to explicit harmful instructions (e.g., ”How to build a bomb?”) through techniques like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), they often fail to address the risks posed by toxic text continuation tasks - where VLMs amplify harmful content when completing partially toxic text prompts paired with semantically adversarial images (Figure 1).
+This vulnerability originates from two research gaps: (a) Current red teaming methods focus on explicit instruction attacks rather than contextual toxicity induction.
+(b) Existing safety protocols neglect cross-modal attack vectors where images amplify output toxicity.
+The multimodal nature of VLMs introduces unique safety challenges: identical text inputs paired with different images can yield drastically different toxicity levels (Liu et al., 2023), sometimes in unintuitive ways.
+For instance, Li et al.
+(Li et al., 2024) demonstrated that pairing jailbreak prompts with even a blank image can significantly increase attack success rates.
+However, systematically identifying images that effectively trigger toxic continuations remains an open and challenging problem.
+Recent progress in white-box jailbreak techniques (Qi et al., 2024; Wang et al., 2024b) has explored the use of adversarial images optimized directly against model parameters.
+While powerful, such gradient-based approaches are impractical in real-world scenarios, particularly within black-box VLM-as-a-Service contexts where full model access is unavailable.
+Additionally, adversarial images generated through gradient-based methods typically exhibit noticeable distortions, making them susceptible to detection and defense mechanisms (Nie et al., 2022; Zhang et al., 2023; Zhao et al., 2024).
+Moreover, studies have demonstrated that these adversarial images typically show limited transferability between VLMs, even after successfully jailbreaking multiple VLMs simultaneously (Schaeffer et al., [n.
+d.]).
+These limitations motivate us to develop stealthy black-box attacks that produce natural adversarial images without full model access.
+Existing black-box jailbreak methods typically adopt suboptimal paradigms, rely on manually engineered pipelines involving typographic manipulations or pre-trained diffusion models to generate query-relevant images (Gong et al., 2023; Liu et al., 2023; Teng et al., 2024).
+However, such methods have critical drawbacks.
+First, structure-based attacks inherently constrain the attack surface, resulting in predictable patterns that are easily mitigated by modern defense mechanisms (Wang et al., 2024a) and detectable by OCR systems.
+Second, such images are primarily designed to bypass sensitive text detection, rather than actively inducing harmful continuations, limiting their applicability to implicit toxicity tasks.
+More recently, automated red-teaming frameworks have emerged to systematically uncover vulnerabilities in VLMs.
+For instance, IDEATOR (Wang et al., 2024c) introduced the ”VLMs against VLMs” paradigm, employing a red-team VLM to generate jailbreak prompts paired with images from general-purpose diffusion models (Rombach et al., 2022).
+While innovative, this approach inherits a fundamental limitation: the inherent conflict between diffusion models’ original training objective (faithful visual-textual alignment) and red teaming goals (toxicity maximization).
+Off-the-shelf diffusers like Stable Diffusion (Rombach et al., 2022) are optimized to avoid generating harmful content during pretraining, creating an intrinsic bias against jailbreaking effectiveness.
+This highlights a critical yet unexplored research opportunity: Can we reconfigure diffusion models to actively induce toxicity while maintaining semantic plausibility?
+Current approaches treat image generation and toxicity induction as decoupled processes - first generating ”safe” images with pretrained diffusers, then hoping they coincidentally trigger harmful outputs.
+Our work overturns this passive paradigm through reinforced diffusion alignment, where the diffusion model is actively tuned to: (1) maximize toxicity through RL-driven reward signals; (2) preserve semantic coherence with image prompts; (3) dynamically adapt to target VLM’s defense patterns.
+By transforming diffusion models from general-purpose generators into active toxicity amplifiers, we establish a new paradigm for multimodal red teaming that fundamentally rethinks how visual modalities can be weaponized against VLMs.
+In this work, we propose Red Team Diffuser (RTD), a novel framework that redefines multimodal jailbreak through semantically adversarial image synthesis.
+Departing from conventional static attack patterns, RTD introduces a dynamic two-phase optimization paradigm.
+First, we leverage the contextual reasoning capabilities of an LLM to perform a greedy search over image prompts, identifying those that maximize the toxicity of VLM responses and are thus highly effective for red teaming.
+Second, the discovered prompts drive RL-based fine-tuning of a diffusion model, using a dual-objective reward: a toxicity reward to amplify the harmfulness of the induced outputs, and an image-prompt alignment reward to ensure semantic consistency as a regularization term.
+This transforms standard diffusion models into toxicity-aware generators, fundamentally altering their output distribution toward vulnerable regions of target VLMs’ multimodal alignment space.
+The main contributions of our work are threefold:
+  • New Attack Paradigm: We propose Red Team Diffuser (RTD), the first framework to integrate reinforcement learning with diffusion models for semantically adversarial image generation, establishing a new paradigm for multimodal red teaming.
+Unlike prior approaches that rely on static patterns or unmodified diffusers, RTD actively co-optimizes both image prompts and diffusers for toxicity induction.
+  • Technical Innovations: We employ an LLM to automatically generate diverse image prompts, followed by iterative toxicity evaluation to identify the most effective triggers.
+Furthermore, we introduce an RL-based fine-tuning strategy with dual rewards, balancing toxicity maximization and semantic preservation, thereby ensuring both high attack potency and natural image quality.
+  • Empirical Contributions: We demonstrate a 10.69% increase in toxicity rate over text-only attacks on LLaVA.
+Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, along with toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s LLaMA.
+Our findings reveal a critical vulnerability in cross-modal toxicity amplification, where semantically adversarial images exacerbate text-based harm.
+2. Related Work
+2.1. Large Vision-Language Models
+Large Vision-Language Models (VLMs) extend Large Language Models (LLMs) by integrating visual modalities through an image encoder and a cross-modal alignment module.
+This design enables VLMs to perform instruction-following and open-ended reasoning over both text and images.
+Recent open-source VLMs have explored various architecture designs.
+LLaVA (Liu et al., 2024a) connects a CLIP encoder (Radford et al., 2021) with LLaMA (Touvron et al., 2023), using GPT-4 (Achiam et al., 2023) to synthesize multimodal instruction-following data for end-to-end fine-tuning.
+MiniGPT-4 (Zhu et al., 2023) employs a linear projection to align a frozen ViT (Dosovitskiy et al., 2020) with Vicuna (Chiang et al., 2023), using pretraining on large-scale image-text pairs followed by high-quality fine-tuning.
+InstructBLIP (Dai et al., 2023), based on BLIP-2 (Li et al., 2023b), introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding.
+While these models achieve impressive performance, the incorporation of visual inputs significantly broadens the attack surface compared to text-only LLMs.
+Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior (Shayegani et al., 2023b; Liu et al., 2024b; Li et al., 2023a; Ma et al., 2025).
+These findings highlight the urgent need to evaluate VLM robustness under multimodal threats—an area this work directly addresses.
+Figure 2.
+Overview of the proposed method.
 Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation.
@@
 Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
-2.2 Jailbreak Attacks against VLMs
+2.2. Jailbreak Attacks against VLMs
 Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content.
 These attacks can be categorized as either white-box or black-box methods.
-White-box jailbreak attacks Bagdasaryan et al.
-(2023); Bailey et al.
-(2023); Shayegani et al.
-(2023a); Carlini et al.
-(2024); Qi et al.
-(2024); Niu et al.
-(2024); Wang et al.
-(2024a) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs.
-While effective, they require full or partial access to model parameters, making them less practical and more easily detectable Nie et al.
-(2022); Zhang et al.
-(2023).
+White-box jailbreak attacks (Bagdasaryan et al., 2023; Bailey et al., 2023; Shayegani et al., 2023a; Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024; Wang et al., 2024b) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs.
+While effective, they require full or partial access to model parameters, making them less practical and more easily detectable (Nie et al., 2022; Zhang et al., 2023).
 In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities.
 Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR.
-For example, FigStep Gong et al.
-(2023) converts harmful text into images using typography to bypass safety mechanisms.
-Similarly, approaches like MM-SafetyBench and VRP Liu et al.
-(2023); Ma et al.
-(2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems.
-To address these limitations, the concept of red team VLMs has advanced with IDEATOR Wang et al.
-(2024b), which autonomously generates jailbreak image-text pairs in a black-box manner.
-By providing a scalable and efficient solution for creating adversarial inputs, IDEATOR has demonstrated the potential of red teaming to expose vulnerabilities in VLMs.
-However, while IDEATOR focuses on leveraging red team VLMs, our work broadens the scope of red teaming by exploring the feasibility of red team diffusers, which are orthogonal to red team VLMs and pose greater challenges.
-Specifically, we investigate the capability of unrestricted images to trigger harmful content generation, highlighting the significant threats posed by such inputs to VLMs.
-Compared to existing approaches, our work presents a more challenging scenario by addressing the limitations of typography-based attacks and provides a rigorous framework for evaluating the robustness of VLMs against adversarial inputs.
-3 Preliminary
-Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm Black et al.
-(2023) to fine-tune a pretrained diffusion model Rombach et al.
-(2022).
-Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning (RL) to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP).
+For example, FigStep (Gong et al., 2023) converts harmful text into images using typography to bypass safety mechanisms.
+Similarly, approaches like MM-SafetyBench and VRP (Liu et al., 2023; Ma et al., 2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems.
+To address these limitations, the concept of red team VLMs has progressed with IDEATOR (Wang et al., 2024c), which autonomously generates jailbreak image-text pairs in a black-box setting.
+By offering a scalable and automated pipeline for discovering jailbreak prompts, IDEATOR demonstrates the promise of using VLMs themselves to expose alignment weaknesses.
+While IDEATOR focuses on prompt-level red teaming, our work introduces a fundamentally different and complementary attack paradigm: reinforced diffusion-based image optimization.
+Rather than relying on VLM-generated inputs, we investigate how semantically adversarial images—crafted via reinforcement learning—can effectively induce toxic text continuations when paired with fixed text prefixes.
+This setup presents a significantly more challenging threat model: it removes reliance on typographic or prompt engineering tricks and explores the vulnerability of VLMs to unrestricted, natural-looking visual stimuli.
+By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks, highlighting a critical blind spot in existing alignment strategies.
+3. Preliminary
+Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm (Black et al., 2023) to fine-tune a pretrained diffusion model (Rombach et al., 2022).
+Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP).
 At each step, actions (denoising predictions) are chosen to maximize the cumulative reward.
@@
 This objective encourages the model to learn a policy that generates samples close to the target distribution.
-To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator Kakade and Langford (2002):
+To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator (Kakade and Langford, 2002):
 $\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg{[}\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg{]},$
@@
 This estimator leverages the ratio of the current model’s likelihood to the previous model’s likelihood, allowing for more efficient updates.
-4 Methodology
-4.1 Threat Model
+4. Methodology
+4.1. Threat Model
 Attack Goals  We focus on the sentence continuation task, an underexplored area in alignment mechanisms that poses unique challenges due to the open-ended nature of model outputs.
@@
 By reinforcement fine-tuning red team diffusers based on the VLM’s feedback, RTD learns the VLM’s behavior patterns and identifies vulnerabilities, refining its policy to trigger toxic content.
-4.2 Greedy Prompt Search
-A central challenge in our task is the design of effective image prompts to trigger harmful content.
-To address this challenge and integrate the toxic continuation task into a multimodal framework, we utilize Google’s Gemini Team et al.
-(2023), disabling its safety settings to generate image prompts for each given sentence.
-Specifically, we instruct Gemini to produce image prompts that preserve the original sentence’s semantics while amplifying harmful themes, thereby increasing the likelihood of generating toxic content.
-Additionally, we leverage in-context learning Brown (2020) by providing examples of previously identified harmful sentence continuations to guide Gemini in generating image prompts consistent with these harmful patterns.
-However, a key limitation arises from the fact that image prompts generated by LLMs may not be optimal for eliciting highly toxic continuations.
-To address this, we propose a greedy prompt search algorithm designed to identify near-optimal image prompts.
-For each sentence, the algorithm iteratively generates image prompts using Gemini and evaluates their associated toxicity scores.
-The toxicity score of each new prompt is compared with the previous one: if the new score is higher, the iteration continues; otherwise, the search terminates.
-4.3 Reward Functions for RTD
-We now introduce the dual-reward mechanism for RTD, which consists of a toxicity reward and an alignment reward.
-Toxicity Reward: In the toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, which is paired with an image $\mathbf{x}$ generated by the RTD.
-The toxicity of the generated continuation is then assessed using the Detoxify classifier Hanu and Unitary team (2020), which evaluates toxicity across six attributes, with scores ranging from 0 (least toxic) to 1 (most toxic).
-The average of these scores is used as the reward function to guide the diffusion model’s optimization process:
+4.2. Greedy Prompt Search
+The effectiveness of our toxic continuation attacks hinges on identifying image prompts that maximally amplify toxicity in VLM-generated outputs.
+To overcome the limitations of directly using LLM-generated prompts, we introduce a dynamic, feedback-driven search procedure designed to discover high-risk image prompts tailored to specific toxic text prefixes.
+We employ Google’s Gemini Pro (Team et al., 2023)—with safety settings disabled—as a red team LLM to propose candidate image prompts conditioned on each text prefix from the LLM safety benchmark.
+Each image prompt is required to retain the semantic context of the prefix while subtly amplifying its harmful intent, thereby increasing the likelihood of inducing toxic completions.
+To enhance control over prompt generation, we apply in-context learning (Brown, 2020) by providing exemplars of previously identified harmful sentence continuations in the system prompt, which serve as implicit guidance for Gemini to produce image prompts consistent with known harmful patterns.
+Despite Gemini’s capabilities, its generated prompts may still be suboptimal in terms of triggering maximal toxicity.
+To this end, we propose a greedy prompt search algorithm to iteratively refine image prompts based on observed toxicity feedback.
+At each iteration, Gemini generates a new candidate prompt, which is used to synthesize an image via the diffusion model.
+This image, paired with the fixed text prefix, is input to the target VLM, and the resulting continuation is scored using a toxicity classifier (e.g., Detoxify (Hanu and Unitary team, 2020)).
+If the toxicity score increases relative to the previous iteration, the new image prompt is retained and the search continues; otherwise, the process terminates.
+This loop yields near-optimal image prompts that serve as strong initialization points for downstream adversarial image generation.
+4.3. Reward Functions for RTD
+A contribution of RTD lies in its dual-reward optimization, which explicitly balances attack effectiveness and semantic plausibility.
+We introduce two complementary rewards: a toxicity reward that drives harmful continuation, and an alignment reward that ensures prompt-consistent image generation.
+Toxicity Reward: In our toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, paired with an image $\mathbf{x}$ generated by the RTD.
+The toxicity of the generated continuation is then assessed using the Detoxify classifier (Hanu and Unitary team, 2020), which evaluates six toxicity-related attributes (e.g., insult, threat, identity attack), each scored between 0 (least toxic) and 1 (most toxic).
+We use the average of these scores as the toxicity reward to guide diffusion-based adversarial image synthesis:
 $\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$
-Alignment Reward: Following Black et al.
-(2023), we introduce an alignment reward based on BERTScore Zhang et al.
-(2019) to improve multimodal understanding.
-In the context of our black-box attack, we expect the generated image $\mathbf{x}$ to be correctly interpreted by the target VLM and aligned with the intended semantics of the given image prompt $p$.
-Furthermore, we treat this alignment reward as a regularization term, helping to mitigate the risk of model collapse.
-Specifically, we define the alignment reward as:
+Alignment Reward: To encourage semantic alignment between the input prompt $p$ and the VLM’s interpretation of the generated image, we introduce an alignment reward that regularizes the image generation process.
+Following (Black et al., 2023), we use BERTScore (Zhang et al., 2019) to measure the similarity between the image prompt and the caption produced by the victim VLM.
+Specifically, the alignment reward is defined as:
 $\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big{(}S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big{)}$
 […]
-Total Reward: The total reward is a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
+Total Reward: The total reward is formulated as a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
 $\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$
-This combined reward allows the RTD to generate both toxic and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence.
+This combined reward allows the RTD to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence.
 The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
@@
 23:  return Fine-tuned Diffuser $D_{\theta}^{*}$.
-5 Experiments
-Table 1: Black-box attack results on the original attack set against LLaVA.
+5. Experiments
+Datasets We use the RealToxicityPrompts benchmark (Gehman et al., 2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2024).
+We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence.
+The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability.
+While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set.
+This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations.
+In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
+Metric: (1) Toxicity Rate.
+To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we use the Detoxify classifier (Hanu and Unitary team, 2020) during training, and both Detoxify and the Perspective API during testing.
+These classifiers assess toxicity across six predefined attributes, with scores ranging from 0 (least toxic) to 1 (most toxic).
+For evaluation, we compute the proportion of continuations whose toxicity score exceeds a threshold of 0.5 for each attribute.
+The use of the Perspective API at test time allows us to assess cross-classifier generalization and mitigate risks of reward hacking or classifier-specific bias.
+(2) CLIPScore.
+To evaluate the semantic alignment between generated images and their prompts, we use CLIPScore (Hessel et al., 2021), a reference-free metric based on the cosine similarity between CLIP-encoded image and text features.
+Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following (Hessel et al., 2021)) for alignment with human judgment.
+In our setting, CLIPScore serves as an auxiliary metric to verify that adversarial images preserve semantic plausibility without compromising visual coherence.
+Implementation Details We use Stable Diffusion v1.5 (Rombach et al., 2022) as the pre-trained diffusion model in our experiments.
+Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM.
+Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation.
+During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU.
+We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
+5.1. Main Results
+To evaluate the effectiveness of Red Team Diffuser (RTD), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM.
+We compare RTD against several baseline strategies to assess its advantage in inducing harmful outputs.
+We first include a Text-only Attack, in which the model is prompted solely with toxic text prefixes without any accompanying image.
+This serves as a unimodal baseline to quantify the additional effect of visual inputs on toxicity.
+Next, we define the Single-Prompt Attack, where Gemini generates a single image prompt per prefix without any optimization.
+This provides a naive multimodal baseline, capturing the effect of directly pairing a non-optimized image with a toxic prefix.
+The Greedy-Prompt Attack improves upon this by iteratively refining the image prompt using our proposed greedy search algorithm.
+Finally, we evaluate the full RTD framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards.
+We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
+Table 1.
+Black-box attack results on the original attack set against LLaVA.
 Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
 (%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
-No Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
-Base Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
-Greedy attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
+Text-only Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
+Single-Prompt Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
+Greedy-Prompt Attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
 RTD w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
@@
 (%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
-No Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
-Base Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
-Greedy attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
+Text-only Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
+Single-Prompt Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
+Greedy-Prompt Attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
 RTD w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
@@
 RTD w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
-Datasets We use the RealToxicityPrompts benchmark Gehman et al.
-(2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations Schick et al.
-(2021); Mehrabi et al.
-(2022); Qi et al.
-(2024).
-We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence.
-The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability.
-While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set.
-This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations.
-In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
-Metric: Toxicity Rate To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark Gehman et al.
-(2020), we utilize the Detoxify classifier Hanu and Unitary team (2020) during training, and both Detoxify and the Perspective API during testing.
-Both classifiers assess toxicity across six predefined attributes.
-Each attribute’s toxicity score ranges from 0 (least toxic) to 1 (most toxic).
-The Perspective API is employed during testing to assess cross-classifier transferability.
-For each attribute, we calculate the proportion of continuations whose toxicity scores exceed a 0.5 threshold.
-Implementation Details We use Stable Diffusion v1.5 Rombach et al.
-(2022) as the pre-trained diffusion model in our experiments.
-Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM.
-Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation.
-During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU.
-We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
-5.1 Main Results
-To evaluate the effectiveness of the Red Team Diffuser (RTD), we conduct black-box attacks on the original attack set and cross-prompt transfer attacks on the hold-out set against LLaVA.
-We compare several attack strategies: Base Attack: Gemini generates a single image prompt without optimization, providing a baseline for comparison.
-Greedy Attack: We iteratively refine the image prompt using the proposed greedy prompt search to maximize toxicity, selecting the most effective prompt at each step.
-RTD: We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
-Table 2: Cross-prompt transfer attack results on the hold-out set against LLaVA.
+Table 2.
+Cross-prompt transfer attack results on the hold-out set against LLaVA.
 Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
 (%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
-No Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
-Base Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
-Greedy attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
+Text-only Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
+Single-Prompt Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
+Greedy-Prompt Attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
 RTD w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
@@
 (%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
-No Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
-Base Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
-Greedy attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
+Text-only Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
+Single-Prompt Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
+Greedy-Prompt Attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
 RTD w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
@@
 RTD w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
-Original Attack Set Table 1 presents the results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API.
-The percentages represent the rate of specific toxic attributes in the outputs, with ‘Any*’ indicating whether the text exhibits at least one of the six toxic attributes.
-The proposed greedy prompt search demonstrates a significant increase in toxicity compared to the Base Attack.
-For example, using the Detoxify classifier, the Greedy Attack achieved an ‘Any*’ rate of 64.12%, compared to 55.72% for the Base Attack.
-This substantial improvement underscores the effectiveness of greedy prompt search in maximizing toxic content generation.
-Furthermore, the Red Team Diffuser (RTD) consistently outperforms the Greedy Attack, with RTD w/ alignment (M) achieving the highest ‘Any*’ rate of 68.70% according to Detoxify.
-This highlights the considerable impact of reinforced fine-tuning in enhancing the RTD’s ability to trigger toxic continuations.
-Although RTD was trained using Detoxify, the results from the Perspective API validate its cross-classifier transferability.
-For example, RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% on Perspective API, demonstrating the robustness of RTD across different toxicity evaluation models.
-Table 3: Cross-model transfer attack results on Gemini-1.5-flash. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
-(%) Detoxify Perspective API
-No Attack 76.23 78.26
-RTD w/o alignment 80.30 81.49
-RTD w/ alignment (L) 79.14 80.83
-RTD w/ alignment (M) 81.33 82.52
-RTD w/ alignment (H) 81.00 82.11
+Original Attack Set Table 1 presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API.
+The percentages represent the rate of specific toxic attributes in the generated outputs, with ‘Any*’ indicating whether a continuation exhibits at least one of the six toxic attributes.
+We begin by examining the Text-only Attack, where only the toxic text prefix is provided without any visual input.
+Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate—achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective.
+This indicates that current alignment mechanisms in VLMs remain insufficient for preventing harmful completions even in unimodal settings, exposing a fundamental vulnerability in their default safety alignment.
+When introducing visual input, the Single-Prompt Attack slightly alters the toxicity pattern but does not significantly improve attack effectiveness, serving as a weak multimodal baseline.
+In contrast, the Greedy-Prompt Attack significantly amplifies toxicity.
+For instance, using the Detoxify classifier, the ‘Any*’ rate increases from 55.72% (Single-Prompt) to 64.12%, illustrating the benefit of greedy prompt search in finding high-risk image prompts.
+The full Red Team Diffuser (RTD) further enhances attack strength.
+RTD w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers.
+Notably, RTD also demonstrates strong cross-classifier transferability: RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify.
+This confirms that the toxic behaviors induced by RTD are not artifacts of a single classifier but reflect a general vulnerability in VLM alignment under multimodal inputs.
+Table 3.
+Cross-model transfer attack results on Gemini-1.5-flash.
+Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
+Detoxify (%)
+(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
+Text-only Attack 4.18 61.75 2.25 34.64 2.84 76.23 76.23
+RTD w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
+RTD w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
+RTD w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
+RTD w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
+Perspective API (%)
+(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
+Text-only Attack 5.76 60.95 3.17 19.14 5.01 74.41 78.26
+RTD w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
+RTD w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
+RTD w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
+RTD w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
 Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA.
-Importantly, the prompts in the hold-out set were not seen during training, making this evaluation a strong indicator of the generalization capability of the proposed methods.
+Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed methods.
 Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set.
 This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM.
-The Greedy Attack demonstrates a clear improvement over the Base Attack according to both the Detoxify classifier and the Perspective API.
-For instance, using Detoxify, the Greedy Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Base Attack.
-Similarly, with the Perspective API, the Greedy Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Base Attack.
-The Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%).
+The Greedy-Prompt Attack demonstrates a clear improvement over the Single-Prompt Attack according to both the Detoxify classifier and the Perspective API.
+For instance, using Detoxify, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Single-Prompt Attack.
+Similarly, with the Perspective API, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Single-Prompt Attack.
+More notably, the full Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%).
 Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%).
-These results clearly validate the effectiveness of our RTD in eliciting toxic continuations from diverse, unseen prompts.
-The consistent performance across both classifiers demonstrates its strong generalization capability beyond the training data.
-Table 4: Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
-(%) Detoxify Perspective API
-No Attack 15.02 18.68
-RTD w/o alignment 41.01 43.90
-RTD w/ alignment (L) 41.85 45.24
-RTD w/ alignment (M) 41.81 44.37
-RTD w/ alignment (H) 41.15 43.78
-5.2 Cross-Model Transferability
-We evaluate the cross-model transferability of the proposed Red Team Diffuser (RTD) by using the RTD trained on LLaVA Liu et al.
-(2024b).
-Transfer attacks are performed on two additional VLMs: Gemini-1.5-flash Team et al.
-(2023) and Llama-3.2-11B-Vision-Instruct.
-Tables 3 and 4 present the ‘Any*’ rates evaluated by both the Detoxify classifier and the Perspective API.
-These results demonstrate that RTD, trained specifically on a single VLM, can effectively trigger toxic content when applied to unseen models, underscoring its robustness and transferability across VLMs.
-Table 3 summarizes the attack performance on Gemini 1.5 Flash.
-RTD w/ alignment (M) achieves the highest ‘Any*’ rates of 81.33% according to Detoxify and 82.52% according to the Perspective API, significantly outperforming other configurations.
-This indicates that the RTD maintains strong transferability to Gemini-1.5-flash, even without further fine-tuning.
-Table 4 presents the results for Llama 3.2 11B Vision-Instruct, a VLM with more advanced alignment mechanisms.
-While the baseline (No Attack) exhibits relatively low toxicity rates (15.02% according to Detoxify and 18.68% according to Perspective API), the RTD significantly increases these metrics.
-Notably, the lower baseline toxicity of Llama 3.2 can be attributed to its high refusal rate for toxic continuation tasks, demonstrating a stronger alignment mechanism that often declines requests to generate potentially harmful content.
-In contrast, other VLMs, such as LLaVA and Gemini, rarely refuse these requests, leading to higher baseline toxicity levels.
-Despite this, RTD is able to effectively bypass Llama 3.2’s refusal mechanisms, significantly increasing toxicity rates.
-Specifically, RTD w/ alignment (L) achieves the highest ‘Any*’ rates of 41.85% according to Detoxify and 45.24% according to Perspective API, closely followed by RTD w/ alignment (M).
-These findings demonstrate that RTD effectively transfers its attack capabilities to VLMs with different architectures and scales, even when those models have stronger alignment mechanisms.
-5.3 Visualizations
-Figure 3 shows a qualitative comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (RTD) (right).
+These consistent gains clearly validate the strong generalization capability of RTD across diverse, unseen prompts and across different toxicity classifiers, underscoring its effectiveness as a black-box multimodal jailbreak method.
+Table 4.
+Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct.
+Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
+Detoxify (%)
+(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
+Text-only Attack 0.25 11.18 0.00 3.92 0.16 14.52 15.02
+RTD w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
+RTD w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
+RTD w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
+RTD w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
+Perspective API (%)
+(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
+Text-only Attack 0.50 13.84 0.00 3.16 0.66 15.59 18.68
+RTD w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
+RTD w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
+RTD w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
+RTD w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
+5.2. Cross-Model Transferability
+We evaluate the cross-model transferability of our proposed method by using the RTD trained solely on LLaVA (Liu et al., 2024a).
+Transfer attacks are performed on two unseen VLMs: Gemini-1.5-flash (Team et al., 2023) and Llama-3.2-11B-Vision-Instruct.
+Tables 3 and 4 present the toxicity rates evaluated by both the Detoxify classifier and the Perspective API.
+These results demonstrate that RTD generalizes effectively across models, highlighting its robustness and potential threat as a transferable multimodal jailbreak method.
+Table 3 presents the results on Gemini-1.5-flash.
+Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective), indicating that harmful completions are already prevalent even without adversarial visual input.
+Furthermore, all RTD variants elevate toxicity levels, with RTD w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively.
+These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images.
+They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios—not just instruction-following tasks.
+Table 4 reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards.
+Here, the Text-only Attack yields substantially lower toxicity rates—15.02% under Detoxify and 18.68% under Perspective API—reflecting a high refusal rate for toxic continuations.
+However, RTD is still able to bypass these safeguards.
+RTD w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective), with RTD (M) performing similarly.
+These results demonstrate RTD’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models.
+Together, they highlight the broader risk of multimodal toxicity attacks against both commercial and open-source VLMs, and underscore the need for more robust, context-aware alignment strategies in future VLM design.
+Figure 3.
+Comparison of images generated by the general-purpose Stable Diffusion (left) and the Red Team Diffuser (right).
+Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
+Table 5. CLIPScore comparison across different diffusers.
+Stable Diffusion RTD (w/o) RTD (L) RTD (M) RTD (H)
+CLIPScore 44.47 41.87 42.86 44.02 44.80
+5.3. Visualizations
+Figure 3 presents a qualitative comparison between images generated by the general-purpose Stable Diffusion model (left) and our Red Team Diffuser (RTD) (right).
 For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2.
-Despite being given the same image prompts, the outputs from the two models differ significantly.
-While the pre-trained Stable Diffusion produces neutral and benign visuals, RTD generates images that subtly evoke discomfort or unease.
+Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications.
+While the general-purpose Stable Diffusion consistently produces neutral and benign visuals.
+In contrast, RTD generates images that subtly evoke discomfort or toxicity-relevant cues.
 For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion.
@@
 Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations.
-The observed differences in visual elements and stylistic features illustrate how RTD effectively adapts image features to exploit vulnerabilities in VLMs, thereby enhancing its ability to induce harmful continuations.
-Figure 3: Comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (right).
-Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
-6 Limitations
-The RTD framework demonstrates strong performance in eliciting harmful outputs from VLMs, but several limitations deserve attention.
-First, this study focuses exclusively on toxic continuation tasks.
-Broader challenges, such as misinformation propagation and ethical biases in complex multimodal scenarios, remain unexplored.
-Second, RTD’s high computational demands, including the use of 8 NVIDIA A100 GPUs, may hinder its adoption by researchers with limited resources.
-Finally, RTD poses ethical risks, as its methods could potentially be misused by adversaries.
-7 Conclusion
-In this work, we propose the Red Team Diffuser (RTD) framework to identify and exploit vulnerabilities in VLMs, specifically in toxic continuation tasks.
-Leveraging greedy prompt search and reinforcement fine-tuning, RTD generates images that effectively trigger toxic continuations, exposing the limitations of existing alignment mechanisms.
-Experimental results validate the effectiveness of RTD, showing significant increases in toxicity rates across both black-box and cross-model settings.
-For instance, RTD raises LLaVA’s toxicity rates by 10.69% on the original attack set and 8.91% on a hold-out set, while also demonstrating strong cross-model transferability, with increases of 5.1% on Gemini and 26.83% on Llama-3.2-11B-Vision-Instruct.
-Qualitative analysis further highlights RTD’s ability to subtly manipulate visual features, generating images that are more provocative and capable of inducing harmful continuations.
-These findings highlight critical gaps in current alignment strategies and emphasize the need for more robust and comprehensive mechanisms to enhance the safety of VLMs in real-world applications.
+The observed differences in visual elements and stylistic features underscore RTD’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs.
+The comparison highlights how reinforced optimization—not merely prompt selection—enables subtle but impactful visual manipulation aligned with toxic continuation goals.
+To complement these qualitative observations, we report CLIPScore (Hessel et al., 2021) as a quantitative measure of image-text semantic alignment.
+As shown in Table 5, all RTD variants achieve CLIPScores comparable to or higher than Stable Diffusion, indicating that the adversarial images remain semantically faithful to their prompts despite conveying more toxic or discomforting signals.
+Notably, RTD with the highest alignment reward ($\lambda=0.2$) attains a CLIPScore of 44.80, surpassing the baseline (44.47) and demonstrating that our stealth-aware reward mechanism enhances both visual regularity and the adversarial effectiveness of the generated images.
+6. Limitations
+While the Red Team Diffuser (RTD) framework shows strong performance and cross-model transferability, it has several limitations: 1.
+Task specificity.
+RTD is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored.
+2.
+Computational cost.
+The RL-based fine-tuning requires significant GPU resources, which may be impractical for resource-constrained users.
+7. Conclusion
+We present Red Team Diffuser (RTD), a novel framework for uncovering critical vulnerabilities in Vision-Language Models (VLMs) through semantically adversarial image generation.
+We formalize cross-modal toxicity amplification as a new attack surface, where visually coherent yet maliciously tuned images exacerbate harmful text continuations—an aspect largely overlooked by current alignment efforts.
+RTD combines greedy prompt search with reinforcement fine-tuning to generate images that reliably trigger toxic outputs.
+Experimental results validate its effectiveness: RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set and 8.91% on a hold-out set, while demonstrating strong cross-model transferability, achieving toxicity gains of 5.1% on Gemini and 26.83% on LLaMA-3.2-11B-Vision-Instruct.
+Qualitative analysis further illustrates how RTD manipulates visual cues to subtly induce toxic continuations.
+These findings expose substantial blind spots in existing VLM alignment strategies and underscore the need for more robust, multimodal-aware safety mechanisms.
+In future work, we aim to expand RTD to broader adversarial objectives and model families, offering deeper insights into VLM vulnerabilities and guiding the development of stronger alignment and defense strategies.

### v4

@@
-Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning
-Abstract.
-The growing deployment of large Vision-Language Models (VLMs) exposes critical safety gaps in their alignment mechanisms.
-While existing jailbreak studies primarily focus on VLMs’ susceptibility to harmful instructions, we reveal a fundamental yet overlooked vulnerability: toxic text continuation – where VLMs produce highly toxic completions when prompted with harmful text prefixes paired with semantically adversarial images.
-To systematically study this threat, we propose Red Team Diffuser (RTD), the first red teaming diffusion model that coordinates adversarial image generation and toxic continuation through reinforcement learning.
-Our key innovations include dynamic cross-modal attack and stealth-aware optimization.
-For toxic text prefixes from an LLM safety benchmark, we conduct greedy search to identify optimal image prompts that maximally induce toxic completions.
-The discovered image prompts then drive RL-based diffusion model fine-tuning, producing semantically-aligned adversarial images that boost toxicity rates.
-Stealth-aware optimization introduces joint adversarial rewards that balance toxicity maximization (via Detoxify classifier) and stealthiness (via BERTScore), circumventing traditional noise-based adversarial patterns.
-Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% over text-only baselines on the original attack set and 8.91% on unseen set, proving generalization capability.
-Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA.
-Our findings expose two critical flaws in current VLM alignment: (1) failure to prevent toxic continuation from harmful prefixes, and (2) overlooking cross-modal attack vectors.
-These results necessitate a paradigm shift toward multimodal red teaming in safety evaluations.
-We will publicly release the RTD framework to facilitate future research on robust VLM alignment.
+RedDiffuser: Red Teaming Vision-Language Models for Toxic Continuation via Reinforced Stable Diffusion
+Abstract
+Large Vision-Language Models (VLMs) remain vulnerable to jailbreak attacks that bypass safety mechanisms and elicit harmful outputs.
+In this work, we examine a subtle yet critical variant of this threat: toxic continuation.
+Unlike standard jailbreaks driven by malicious instructions, toxic continuation arises when VLMs are exposed to harmful prefixes paired with adversarial images, leading to toxic completions.
+To address this, we propose RedDiffuser (RedDiff), the first multimodal red teaming framework that leverages reinforcement learning to fine-tune diffusion models for generating natural-looking adversarial images.
+RedDiffuser combines greedy search to identify image prompts that maximize toxicity with reinforcement optimization that balances two objectives: (1) promoting toxic continuation, guided by a toxicity classifier, and (2) maintaining semantic coherence and stealthiness, measured by text–image similarity scores.
+This design circumvents traditional noise-based adversarial patterns and yields more realistic attack vectors.
+Experiments show that RedDiffuser significantly increases the toxicity rate of LLaVA-1.5 outputs by 10.69% and 8.91% on the original and hold-out sets, respectively.
+Moreover, it demonstrates strong transferability, raising toxicity rates by 5.1% on Gemini-1.5-flash and 26.83% on Llama-3.2-11B-Vision-Instruct.
+We also consider a more realistic deployment setting involving external NSFW guardrails.
+Incorporating a checker-aware training signal allows RedDiffuser to retain attack effectiveness while achieving high guardrail pass rates, indicating that toxic continuation vulnerabilities persist under safety-filtered conditions.
+These findings uncover two overlooked flaws in current VLM alignment: the failure to mitigate toxic continuation and the vulnerability to cross-modal adversarial amplification.
+Our results highlight the urgency of multimodal red teaming in evaluating VLM safety.
+We will release the RedDiffuser codebase to support future research.
 Disclaimer: This paper contains content that may be disturbing or offensive.
-Red Teaming, Vision-Language Models, Diffusion Models, Reinforcement Learning
-ccs: Computing methodologies Artificial intelligence
-1. Introduction
-Figure 1.
-Example of a successful jailbreak on LLaVA.
-The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as general-purpose diffusion model (Rombach et al., 2022).
-The rapid deployment of large Vision-Language Models (VLMs), such as GPT-4o (Achiam et al., 2023) and Gemini (Team et al., 2023), has exposed critical blind spots in their multimodal alignment mechanisms.
-While existing safety strategies predominantly focus on preventing responses to explicit harmful instructions (e.g., ”How to build a bomb?”) through techniques like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), they often fail to address the risks posed by toxic text continuation tasks - where VLMs amplify harmful content when completing partially toxic text prompts paired with semantically adversarial images (Figure 1).
-This vulnerability originates from two research gaps: (a) Current red teaming methods focus on explicit instruction attacks rather than contextual toxicity induction.
-(b) Existing safety protocols neglect cross-modal attack vectors where images amplify output toxicity.
-The multimodal nature of VLMs introduces unique safety challenges: identical text inputs paired with different images can yield drastically different toxicity levels (Liu et al., 2023), sometimes in unintuitive ways.
-For instance, Li et al.
-(Li et al., 2024) demonstrated that pairing jailbreak prompts with even a blank image can significantly increase attack success rates.
-However, systematically identifying images that effectively trigger toxic continuations remains an open and challenging problem.
-Recent progress in white-box jailbreak techniques (Qi et al., 2024; Wang et al., 2024b) has explored the use of adversarial images optimized directly against model parameters.
-While powerful, such gradient-based approaches are impractical in real-world scenarios, particularly within black-box VLM-as-a-Service contexts where full model access is unavailable.
-Additionally, adversarial images generated through gradient-based methods typically exhibit noticeable distortions, making them susceptible to detection and defense mechanisms (Nie et al., 2022; Zhang et al., 2023; Zhao et al., 2024).
-Moreover, studies have demonstrated that these adversarial images typically show limited transferability between VLMs, even after successfully jailbreaking multiple VLMs simultaneously (Schaeffer et al., [n.
-d.]).
-These limitations motivate us to develop stealthy black-box attacks that produce natural adversarial images without full model access.
-Existing black-box jailbreak methods typically adopt suboptimal paradigms, rely on manually engineered pipelines involving typographic manipulations or pre-trained diffusion models to generate query-relevant images (Gong et al., 2023; Liu et al., 2023; Teng et al., 2024).
-However, such methods have critical drawbacks.
-First, structure-based attacks inherently constrain the attack surface, resulting in predictable patterns that are easily mitigated by modern defense mechanisms (Wang et al., 2024a) and detectable by OCR systems.
-Second, such images are primarily designed to bypass sensitive text detection, rather than actively inducing harmful continuations, limiting their applicability to implicit toxicity tasks.
-More recently, automated red-teaming frameworks have emerged to systematically uncover vulnerabilities in VLMs.
-For instance, IDEATOR (Wang et al., 2024c) introduced the ”VLMs against VLMs” paradigm, employing a red-team VLM to generate jailbreak prompts paired with images from general-purpose diffusion models (Rombach et al., 2022).
-While innovative, this approach inherits a fundamental limitation: the inherent conflict between diffusion models’ original training objective (faithful visual-textual alignment) and red teaming goals (toxicity maximization).
-Off-the-shelf diffusers like Stable Diffusion (Rombach et al., 2022) are optimized to avoid generating harmful content during pretraining, creating an intrinsic bias against jailbreaking effectiveness.
-This highlights a critical yet unexplored research opportunity: Can we reconfigure diffusion models to actively induce toxicity while maintaining semantic plausibility?
-Current approaches treat image generation and toxicity induction as decoupled processes - first generating ”safe” images with pretrained diffusers, then hoping they coincidentally trigger harmful outputs.
-Our work overturns this passive paradigm through reinforced diffusion alignment, where the diffusion model is actively tuned to: (1) maximize toxicity through RL-driven reward signals; (2) preserve semantic coherence with image prompts; (3) dynamically adapt to target VLM’s defense patterns.
-By transforming diffusion models from general-purpose generators into active toxicity amplifiers, we establish a new paradigm for multimodal red teaming that fundamentally rethinks how visual modalities can be weaponized against VLMs.
-In this work, we propose Red Team Diffuser (RTD), a novel framework that redefines multimodal jailbreak through semantically adversarial image synthesis.
-Departing from conventional static attack patterns, RTD introduces a dynamic two-phase optimization paradigm.
-First, we leverage the contextual reasoning capabilities of an LLM to perform a greedy search over image prompts, identifying those that maximize the toxicity of VLM responses and are thus highly effective for red teaming.
-Second, the discovered prompts drive RL-based fine-tuning of a diffusion model, using a dual-objective reward: a toxicity reward to amplify the harmfulness of the induced outputs, and an image-prompt alignment reward to ensure semantic consistency as a regularization term.
-This transforms standard diffusion models into toxicity-aware generators, fundamentally altering their output distribution toward vulnerable regions of target VLMs’ multimodal alignment space.
-The main contributions of our work are threefold:
-  • New Attack Paradigm: We propose Red Team Diffuser (RTD), the first framework to integrate reinforcement learning with diffusion models for semantically adversarial image generation, establishing a new paradigm for multimodal red teaming.
-Unlike prior approaches that rely on static patterns or unmodified diffusers, RTD actively co-optimizes both image prompts and diffusers for toxicity induction.
-  • Technical Innovations: We employ an LLM to automatically generate diverse image prompts, followed by iterative toxicity evaluation to identify the most effective triggers.
-Furthermore, we introduce an RL-based fine-tuning strategy with dual rewards, balancing toxicity maximization and semantic preservation, thereby ensuring both high attack potency and natural image quality.
-  • Empirical Contributions: We demonstrate a 10.69% increase in toxicity rate over text-only attacks on LLaVA.
-Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, along with toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s LLaMA.
-Our findings reveal a critical vulnerability in cross-modal toxicity amplification, where semantically adversarial images exacerbate text-based harm.
-2. Related Work
-2.1. Large Vision-Language Models
+I Introduction
+The rapid deployment of Vision-Language Models (VLMs) such as GPT-4o [achiam2023gpt] and Gemini [team2023gemini] has highlighted significant vulnerabilities in their safety alignment.
+Recent research demonstrates that VLMs are susceptible to jailbreak attacks, where adversaries can bypass built-in safety mechanisms and elicit harmful outputs.
+Among these, a more insidious threat emerges: toxic continuation.
+Despite progress in refusing explicit harmful instructions, existing defenses fall short in addressing this specific vulnerability, where the model, when presented with both a malicious input and an initial segment of harmful output, can be manipulated to generate highly toxic completions.
+This vulnerability is particularly pronounced in the multimodal setting, where images can substantially amplify the risk of toxic continuation (Figure 1).
+These gaps highlight the urgent need for systematic investigation and mitigation of this emerging risk, moving beyond the conventional focus on single-modality and instruction-based jailbreaks.
+However, a critical challenge in addressing this multimodal vulnerability is identifying the very images that induce such toxic responses, a problem that remains largely unsolved.
+Li et al.
+[li2024images] show that even a blank image can sharply increase the success rate of jailbreak attacks, demonstrating that the connection between image content and VLM toxicity can defy intuitive understanding.
+While some white-box approaches employ gradient-based optimization to generate adversarial images [qi2024visual, wang2024white], these methods typically require full access to the model and often result in images with visible artifacts that are unnatural and easily detected [nie2022diffusion, zhang2023mutation].
+Furthermore, their limited transferability across different VLMs [schaefferfailures] severely restricts their real-world applicability.
+Figure 1: Example of toxic continuation on LLaVA. RedDiffuser generates images capable of inducing harmful continuations, whereas images from a standard diffusion model lead to benign VLM output.
+These challenges highlight the need for stealthy, black-box attacks capable of generating natural-looking yet adversarial images without requiring access to model internals.
+Most existing black-box jailbreaks depend on hand-crafted pipelines or unmodified diffusion models [gong2023figstep, liu2023query], which constrain the range of possible attacks and focus mainly on evading text-based filters rather than proactively inducing toxic model responses.
+Recent methods such as IDEATOR [wang2024ideator] present a “VLMs against VLMs” approach, pairing prompts with images generated by off-the-shelf diffusion models.
+However, these diffusion models are typically pretrained to improve aesthetic quality [rombach2022high], making them inherently misaligned with the objectives of red teaming.
+This gap leads to a fundamental question: Can we directly optimize diffusion models to actively induce toxic continuations in VLMs while preserving image naturalness?
+We address this challenge by introducing RedDiffuser (RedDiff), the first framework to employ reinforcement learning for red teaming with diffusion models.
+RedDiffuser operates in two coordinated phases.
+In the first phase, an LLM conducts a greedy search over candidate image prompts to identify those that maximize toxicity in the VLM’s output.
+These selected prompts then serve as the basis for the second phase, where the diffusion model is reinforcement fine-tuned with a dual reward: one term encourages toxic output (measured by Detoxify), while the other promotes semantic coherence (measured by BERTScore).
+This process turns the diffusion model from a neutral generator into an adversarial tool for exposing alignment vulnerabilities in VLMs.
+We also extend RedDiffuser to handle scenarios where external NSFW guardrails are present, introducing a checker-aware variant.
+Our main contributions are as follows:
+  • We present RedDiffuser, the first framework to integrate reinforcement learning with diffusion models for generating semantically adversarial images, thereby enabling effective cross-modal jailbreaks.
+  • Our approach combines greedy prompt search with reinforcement fine-tuning, guided by both toxicity and alignment rewards, to ensure strong attack efficacy while maintaining natural image quality.
+  • RedDiffuser increases toxicity rates on LLaVA by 10.69% over text-only attacks, achieves 8.91% on a hold-out set, and transfers successfully to Gemini (+5.1%) and LLaMA-Vision (+26.83%).
+  • Under external NSFW guardrails, the checker-aware variant of RedDiffuser improves the guardrail pass rate by +40.4% over Stable Diffusion, indicating that the attack remains feasible despite safety filtering.
+II Related Work
+II-A Vision-Language Models
 Large Vision-Language Models (VLMs) extend Large Language Models (LLMs) by integrating visual modalities through an image encoder and a cross-modal alignment module.
@@
 Recent open-source VLMs have explored various architecture designs.
-LLaVA (Liu et al., 2024a) connects a CLIP encoder (Radford et al., 2021) with LLaMA (Touvron et al., 2023), using GPT-4 (Achiam et al., 2023) to synthesize multimodal instruction-following data for end-to-end fine-tuning.
-MiniGPT-4 (Zhu et al., 2023) employs a linear projection to align a frozen ViT (Dosovitskiy et al., 2020) with Vicuna (Chiang et al., 2023), using pretraining on large-scale image-text pairs followed by high-quality fine-tuning.
-InstructBLIP (Dai et al., 2023), based on BLIP-2 (Li et al., 2023b), introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding.
+LLaVA [liu2024visual] connects a CLIP encoder [radford2021learning] with LLaMA [touvron2023llama], using GPT-4 [achiam2023gpt] to synthesize multimodal instruction-following data for end-to-end fine-tuning.
+MiniGPT-4 [zhu2023minigpt] employs a linear projection to align a frozen ViT [dosovitskiy2020image] with Vicuna [chiang2023vicuna], using pretraining on large-scale image-text pairs followed by high-quality fine-tuning.
+InstructBLIP [dai2023instructblip], based on BLIP-2 [li2023blip], introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding.
 While these models achieve impressive performance, the incorporation of visual inputs significantly broadens the attack surface compared to text-only LLMs.
-Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior (Shayegani et al., 2023b; Liu et al., 2024b; Li et al., 2023a; Ma et al., 2025).
+Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior [shayegani2023survey, liu2024survey, li2023privacy, ma2025safety].
 These findings highlight the urgent need to evaluate VLM robustness under multimodal threats—an area this work directly addresses.
-Figure 2.
-Overview of the proposed method.
-Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation.
-The diffusion model generates an image based on this prompt.
-Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
-2.2. Jailbreak Attacks against VLMs
+II-B Jailbreak Attacks on VLMs
 Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content.
 These attacks can be categorized as either white-box or black-box methods.
-White-box jailbreak attacks (Bagdasaryan et al., 2023; Bailey et al., 2023; Shayegani et al., 2023a; Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024; Wang et al., 2024b) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs.
-While effective, they require full or partial access to model parameters, making them less practical and more easily detectable (Nie et al., 2022; Zhang et al., 2023).
-In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities.
-Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR.
-For example, FigStep (Gong et al., 2023) converts harmful text into images using typography to bypass safety mechanisms.
-Similarly, approaches like MM-SafetyBench and VRP (Liu et al., 2023; Ma et al., 2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems.
-To address these limitations, the concept of red team VLMs has progressed with IDEATOR (Wang et al., 2024c), which autonomously generates jailbreak image-text pairs in a black-box setting.
+White-box jailbreak attacks [bagdasaryan2023ab, bailey2023image, shayegani2023jailbreak, carlini2024aligned, qi2024visual, niu2024jailbreaking, wang2024white] typically leverage gradient information to perturb input images or text, thereby increasing the likelihood of harmful outputs.
+While effective, they require full or partial access to model parameters, which makes them less practical and easier to detect [nie2022diffusion, zhang2023mutation].
+In contrast, black-box jailbreak attacks provide more practical and scalable solutions because they do not require direct access to model parameters but instead exploit external vulnerabilities.
+Despite these advantages, many existing black-box approaches rely heavily on typography-based manipulations, which are vulnerable to detection techniques such as OCR.
+For example, FigStep [gong2023figstep] converts harmful text into images using typography to bypass safety mechanisms.
+Similarly, approaches like MM-SafetyBench and VRP [liu2023query, ma2024visual] also incorporate typography-driven attacks, reducing their stealthiness against advanced detection systems.
+To overcome these limitations, the concept of red team VLMs has advanced with IDEATOR [wang2024ideator], which autonomously generates jailbreak image-text pairs in a black-box setting.
 By offering a scalable and automated pipeline for discovering jailbreak prompts, IDEATOR demonstrates the promise of using VLMs themselves to expose alignment weaknesses.
 While IDEATOR focuses on prompt-level red teaming, our work introduces a fundamentally different and complementary attack paradigm: reinforced diffusion-based image optimization.
-Rather than relying on VLM-generated inputs, we investigate how semantically adversarial images—crafted via reinforcement learning—can effectively induce toxic text continuations when paired with fixed text prefixes.
-This setup presents a significantly more challenging threat model: it removes reliance on typographic or prompt engineering tricks and explores the vulnerability of VLMs to unrestricted, natural-looking visual stimuli.
-By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks, highlighting a critical blind spot in existing alignment strategies.
-3. Preliminary
-Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm (Black et al., 2023) to fine-tune a pretrained diffusion model (Rombach et al., 2022).
-Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP).
-At each step, actions (denoising predictions) are chosen to maximize the cumulative reward.
-The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps.
-Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$.
-The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process.
-The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$.
-The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$.
-That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
+Instead of relying on VLM-generated inputs, we investigate how semantically adversarial images crafted via reinforcement learning can induce toxic text continuations when paired with fixed text prefixes.
+This setup constitutes a significantly more challenging threat model, as it removes dependence on typographic or prompt engineering tricks and examines the vulnerability of VLMs to unrestricted, natural-looking visual stimuli.
+By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks and highlight a critical blind spot in existing alignment strategies.
+Figure 2: RedDiffuser overview.
+Given an incomplete toxic sentence, Gemini selects an image prompt via greedy search.
+A diffusion model generates an image, which is passed to a VLM (LLaVA) to produce a continuation.
+Toxicity and alignment scores from Detoxify and BERTScore are used as rewards to fine-tune the diffusion model.
+III Preliminary
+[…]
 The goal of DDPO is to maximize the expected cumulative reward, which is defined over the sample distribution $p_{\theta}(x_{0}\mid c)$ and the context distribution $p(c)$:
-$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right].$
+$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right]$ (1)
 This objective encourages the model to learn a policy that generates samples close to the target distribution.
-To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator (Kakade and Langford, 2002):
-$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg{[}\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg{]},$
+To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator [kakade2002approximately]:
+$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg[\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg]$ (2)
 where the expectation is taken over denoising trajectories generated by the previous policy $\pi_{\theta_{\text{old}}}$.
 This estimator leverages the ratio of the current model’s likelihood to the previous model’s likelihood, allowing for more efficient updates.
-4. Methodology
-4.1. Threat Model
-Attack Goals  We focus on the sentence continuation task, an underexplored area in alignment mechanisms that poses unique challenges due to the open-ended nature of model outputs.
-The attacker aims to bypass the model’s safety measures and induce toxic continuations, such as the generation of obscene or profane language.
-Toxic continuations can also be viewed as a form of multimodal prompt injection, where carefully designed text prefixes and image triggers manipulate the model’s output into generating harmful content.
-Adversary Capabilities  We assume the attacker has only black-box access to the victim VLM, simulating real-world interactions via external interfaces.
-By reinforcement fine-tuning red team diffusers based on the VLM’s feedback, RTD learns the VLM’s behavior patterns and identifies vulnerabilities, refining its policy to trigger toxic content.
-4.2. Greedy Prompt Search
-The effectiveness of our toxic continuation attacks hinges on identifying image prompts that maximally amplify toxicity in VLM-generated outputs.
-To overcome the limitations of directly using LLM-generated prompts, we introduce a dynamic, feedback-driven search procedure designed to discover high-risk image prompts tailored to specific toxic text prefixes.
-We employ Google’s Gemini Pro (Team et al., 2023)—with safety settings disabled—as a red team LLM to propose candidate image prompts conditioned on each text prefix from the LLM safety benchmark.
-Each image prompt is required to retain the semantic context of the prefix while subtly amplifying its harmful intent, thereby increasing the likelihood of inducing toxic completions.
-To enhance control over prompt generation, we apply in-context learning (Brown, 2020) by providing exemplars of previously identified harmful sentence continuations in the system prompt, which serve as implicit guidance for Gemini to produce image prompts consistent with known harmful patterns.
-Despite Gemini’s capabilities, its generated prompts may still be suboptimal in terms of triggering maximal toxicity.
-To this end, we propose a greedy prompt search algorithm to iteratively refine image prompts based on observed toxicity feedback.
-At each iteration, Gemini generates a new candidate prompt, which is used to synthesize an image via the diffusion model.
-This image, paired with the fixed text prefix, is input to the target VLM, and the resulting continuation is scored using a toxicity classifier (e.g., Detoxify (Hanu and Unitary team, 2020)).
-If the toxicity score increases relative to the previous iteration, the new image prompt is retained and the search continues; otherwise, the process terminates.
-This loop yields near-optimal image prompts that serve as strong initialization points for downstream adversarial image generation.
-4.3. Reward Functions for RTD
-A contribution of RTD lies in its dual-reward optimization, which explicitly balances attack effectiveness and semantic plausibility.
-We introduce two complementary rewards: a toxicity reward that drives harmful continuation, and an alignment reward that ensures prompt-consistent image generation.
-Toxicity Reward: In our toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, paired with an image $\mathbf{x}$ generated by the RTD.
-The toxicity of the generated continuation is then assessed using the Detoxify classifier (Hanu and Unitary team, 2020), which evaluates six toxicity-related attributes (e.g., insult, threat, identity attack), each scored between 0 (least toxic) and 1 (most toxic).
-We use the average of these scores as the toxicity reward to guide diffusion-based adversarial image synthesis:
-$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$
-Alignment Reward: To encourage semantic alignment between the input prompt $p$ and the VLM’s interpretation of the generated image, we introduce an alignment reward that regularizes the image generation process.
-Following (Black et al., 2023), we use BERTScore (Zhang et al., 2019) to measure the similarity between the image prompt and the caption produced by the victim VLM.
-Specifically, the alignment reward is defined as:
-$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big{(}S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big{)}$
-[…]
+IV Methodology
+IV-A Threat Model
+Attack Goals.
+We target the sentence continuation task, where the model generates open-ended completions from partial inputs.
+The attacker’s objective is to bypass safety mechanisms and induce toxic content, including profanity, insults, or threats, by pairing adversarial images with fixed text prefixes.
+This constitutes a form of multimodal prompt injection, where carefully crafted image-text inputs bias the model toward unsafe responses.
+Adversary Capabilities.
+We consider a black-box setting where the attacker interacts with the VLM solely through input-output queries, as in real-world API access.
+By fine-tuning the diffusion model via reinforcement learning with feedback from the VLM, the adversary learns to exploit alignment weaknesses and increase the likelihood of toxic completions.
+IV-B Greedy Prompt Search
+Toxic continuation attacks require identifying image prompts that can consistently amplify toxicity in VLM outputs.
+However, directly using prompts generated by a general LLM is often suboptimal.
+To overcome this limitation, we design a feedback‑driven greedy prompt search procedure that iteratively refines image prompts to maximize toxicity amplification for given toxic text prefixes.
+In our approach, we employ Google’s Gemini Pro [team2023gemini] with safety settings disabled as the red‑team LLM to propose candidate image prompts.
+To guide Gemini’s output toward semantically consistent yet harmful prompts, we utilize in-context learning [brown2020language] by providing Gemini with exemplars of previously identified harmful continuations.
+The greedy prompt search refines these initial prompts iteratively using feedback from the target VLM and a toxicity classifier (e.g., Detoxify [Detoxify]).
+At each iteration, Gemini proposes a new candidate prompt, from which the diffusion model synthesizes a semantically adversarial image.
+The toxicity of the resulting VLM continuation is then evaluated, given the generated image and fixed text prefix.
+If the toxicity score improves over the previous iteration, the new prompt is retained; otherwise, the process terminates.
+The resulting optimized prompts provide strong initialization points for subsequent reinforcement fine-tuning of the diffusion model.
+IV-C Reward Functions for RedDiffuser
+After identifying adversarial image prompts through greedy search, we further fine-tune the diffusion model using reinforcement learning guided by a dual reward: a toxicity reward, which encourages harmful continuations, and an alignment reward, which maintains semantic consistency between the image and its image prompt.
+Toxicity Reward: Given a fixed text prefix $\mathbf{y}$ and an image $\mathbf{x}$ generated by RedDiffuser, we prompt the target VLM $\mathcal{M}$ to complete the sentence.
+The continuation is scored by Detoxify [Detoxify], which evaluates six toxicity attributes (e.g., insult, threat).
+We use the average score as the toxicity reward:
+$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$ (3)
+Alignment Reward: To ensure that generated images remain faithful to their image prompts, we introduce an alignment reward inspired by [black2023training].
+Specifically, we use BERTScore [zhang2019bertscore] to measure semantic similarity between the image prompt $p$ and the VLM’s interpretation of the generated image:
+$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big(S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big)$ (4)
+where $\mathcal{M}(\mathbf{x},\text{``Describe the image in short."})$ is the generated image description from the VLM.
+$S$ calculates the cosine similarity of BERT embeddings between $p_{i}$ and $\mathcal{M}(\mathbf{x},\text{``Describe..."}))_{j}$.
+$|p|$ is the number of words in the image prompt.
+This alignment reward acts as a semantic constraint to avoid prompt-irrelevant image drift and model collapse during optimization.
 Total Reward: The total reward is formulated as a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
-$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$
-This combined reward allows the RTD to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence.
-The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
-Algorithm 1 Training Procedure of Red Team Diffuser
+$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$ (5)
+This combined reward allows the RedDiffuser to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence.
+The overview of RedDiffuser is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
+Algorithm 1 Training Procedure of RedDiffuser
 Require: Pre-trained Diffuser $D_{\theta}$, frozen LLM (e.g., Gemini) $\mathcal{G}$, target VLM $\mathcal{M}$, sentence continuation corpus $Y:=\{\mathbf{y}^{i}\}_{i=1}^{m}$, batch size $b$, learning rate $\alpha$, policy update steps $N$;
@@
 14:end for
-15:  for $step=1,...,N$ do
+15:for step=1,,Nstep=1,...,N do
 16:  Sample a batch of sentence-prompt pairs $\{(\mathbf{y}^{k},p^{k})\}_{k=1}^{b}$ from $Y$ and $P$.
@@
 22:end for
-23:  return Fine-tuned Diffuser $D_{\theta}^{*}$.
-5. Experiments
-Datasets We use the RealToxicityPrompts benchmark (Gehman et al., 2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2024).
+23:return Fine-tuned RedDiffuser $D_{\theta}^{*}$.
+IV-D Extensions for Attacking VLMs with External Guardrails
+Real-world deployments often place external guardrails (e.g., image-level NSFW filters) upstream of VLMs, which can block visually explicit images before they influence downstream text generation.
+To evaluate attacks under this operational constraint, we consider a setting where generated images must both induce toxic continuations and be accepted by external checkers.
+Concretely, we adopt a dual-guardrail setup.
+The first guardrail is the Stable Diffusion image-level NSFW checker; the second is a VLM-based semantic checker implemented by prompting a VLM to judge whether an image contains sexual, violent, or otherwise NSFW content.
+We treat an image as safe only if both checkers classify it as non-NSFW; images flagged by either checker are refused and do not produce continuations.
+To adapt RedDiffuser to this setting, we introduce two complementary extensions: (1) light-weight modifications to the greedy prompt search that incorporate guardrail signals to re-score candidate prompts, and (2) a checker-aware reward shaping mechanism that masks the toxicity reward when a generated image is flagged by the external safety checker.
+Greedy Prompt Search with Guardrail Signals
+Given the image prompts obtained from the base greedy search, we generate ten images per prompt due to the stochasticity of diffusion sampling.
+If any sampled image passes both guardrails, the prompt is retained.
+If all attempts are blocked, we regenerate the image prompt by prompting the attacker LLM with explicit awareness of the external guardrails; the regeneration then follows the same greedy search procedure as before, except that any image prompt that produces NSFW images is skipped.
+This selective filter-and-regenerate step minimally adapts the prompt set to the guardrail-augmented setting.
+Checker-Aware Reward Shaping
+When external guardrails are applied, images that fail either check are directly refused.
+To account for this, we introduce a binary indicator function $\mathbb{I}(\mathbf{x})$ that determines whether a generated image $\mathbf{x}$ passes the dual guardrails:
+$\mathbb{I}(\mathbf{x})=\mathbf{1}\{g_{1}(\mathbf{x})=\text{safe}\ \land\ g_{2}(\mathbf{x})=\text{safe}\}$ (6)
+where $\mathbf{1}\{\cdot\}$ denotes the indicator function, and $g_{1}$ and $g_{2}$ denote the Stable Diffusion NSFW checker and the VLM-based semantic checker, respectively.
+We condition the toxicity reward on this indicator:
+$\mathcal{R}_{\text{tox}}^{\text{guard}}(\mathbf{x},\mathbf{y})=\mathbb{I}(\mathbf{x})\cdot\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y})).$ (7)
+The alignment reward remains unchanged, giving the total reward:
+$\mathcal{R}_{\text{total}}^{\text{guard}}=\mathcal{R}_{\text{tox}}^{\text{guard}}+\lambda\mathcal{R}_{\text{align}}$ (8)
+This checker-aware shaping drives RedDiffuser to discover adversarial image styles that are subtle enough to evade guardrails while remaining effective at eliciting toxic continuations.
+V Experiments
+Datasets We use the RealToxicityPrompts benchmark [gehman2020realtoxicityprompts], specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations [schick2021self, mehrabi2022robust, qi2024visual].
 We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence.
@@
 While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set.
-This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations.
-In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
-Metric: (1) Toxicity Rate.
-To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we use the Detoxify classifier (Hanu and Unitary team, 2020) during training, and both Detoxify and the Perspective API during testing.
+This characteristic makes them a more rigorous benchmark for evaluating RedDiffuser’s ability to induce harmful continuations.
+In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RedDiffuser’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
+Metrics: (1) Toxicity Rate.
+To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark [gehman2020realtoxicityprompts], we use the Detoxify classifier [Detoxify] during training, and both Detoxify and the Perspective API during testing.
 These classifiers assess toxicity across six predefined attributes, with scores ranging from 0 (least toxic) to 1 (most toxic).
 For evaluation, we compute the proportion of continuations whose toxicity score exceeds a threshold of 0.5 for each attribute.
-The use of the Perspective API at test time allows us to assess cross-classifier generalization and mitigate risks of reward hacking or classifier-specific bias.
+The use of the Perspective API at test time enables cross-classifier validation, helping to verify that the RedDiffuser’s behavior is not overfitted to Detoxify and thus not a result of reward hacking or classifier-specific bias.
 (2) CLIPScore.
-To evaluate the semantic alignment between generated images and their prompts, we use CLIPScore (Hessel et al., 2021), a reference-free metric based on the cosine similarity between CLIP-encoded image and text features.
-Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following (Hessel et al., 2021)) for alignment with human judgment.
+To evaluate the semantic alignment between the generated images and the image prompts, we use CLIPScore [hessel2021clipscore], a reference-free metric based on the cosine similarity between CLIP-encoded image and text features.
+Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following [hessel2021clipscore]) for alignment with human judgment.
 In our setting, CLIPScore serves as an auxiliary metric to verify that adversarial images preserve semantic plausibility without compromising visual coherence.
-Implementation Details We use Stable Diffusion v1.5 (Rombach et al., 2022) as the pre-trained diffusion model in our experiments.
+(3) Guardrail Pass Rate (GPR).
+When external image-level guardrails are applied, images flagged as NSFW are blocked before producing a continuation.
+To measure how often generated images remain admissible, we report the Guardrail Pass Rate (GPR), defined as the average value of the indicator in Eq.
+(6):
+$\text{GPR}=\frac{1}{N}\sum_{\mathbf{x}}\mathbb{I}(\mathbf{x})$ (9)
+where $N$ is the number of samples and $\mathbb{I}(\mathbf{x})$ indicates whether the image passes both guardrails.
+Implementation Details We use Stable Diffusion v1.5 [rombach2022high] as the pre-trained diffusion model in our experiments.
 Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM.
-Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation.
-During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU.
-We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
-5.1. Main Results
-To evaluate the effectiveness of Red Team Diffuser (RTD), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM.
-We compare RTD against several baseline strategies to assess its advantage in inducing harmful outputs.
+Fine-tuning RedDiffuser is conducted using 8 NVIDIA A100 GPUs, with 6 for training and 2 for reward computation.
+Reinforcement fine-tuning uses a batch size of 24 per GPU and a learning rate of 3e-4.
+We apply early stopping after 600 updates to mitigate reward hacking [gao2023scaling].
+We experiment with four alignment reward weights $\lambda$: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
+TABLE I: Black-box attack results on the original attack set against LLaVA.
+Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
+Any* indicates the text exhibits at least one of the six toxic attributes.
+Detoxify (%)
+(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
+Text-Only Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
+Single-Prompt Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
+Greedy-Prompt Attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
+RedDiff w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
+RedDiff w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
+RedDiff w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
+RedDiff w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
+Perspective API (%)
+(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
+Text-Only Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
+Single-Prompt Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
+Greedy-Prompt Attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
+RedDiff w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
+RedDiff w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
+RedDiff w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
+RedDiff w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
+V-A Main Results
+To evaluate the effectiveness of RedDiffuser (RedDiff), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM.
+We compare RedDiffuser against several baseline strategies to assess its advantage in inducing harmful outputs.
 We first include a Text-Only Attack, in which the model is prompted solely with toxic text prefixes without any accompanying image.
@@
 The Greedy-Prompt Attack improves upon this by iteratively refining the image prompt using our proposed greedy search algorithm.
-Finally, we evaluate the full RTD framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards.
-We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
-Table 1.
-Black-box attack results on the original attack set against LLaVA.
-Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
-‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
-Detoxify (%)
-(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
-Text-only Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
-Single-Prompt Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
-Greedy-Prompt Attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
-RTD w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
-RTD w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
-RTD w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
-RTD w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
-Perspective API (%)
-(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
-Text-only Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
-Single-Prompt Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
-Greedy-Prompt Attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
-RTD w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
-RTD w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
-RTD w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
-RTD w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
-Table 2.
-Cross-prompt transfer attack results on the hold-out set against LLaVA.
+Finally, we evaluate the full RedDiffuser framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards.
+We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RedDiff w/ alignment (L)), 0.1 (RedDiff w/ alignment (M)), and 0.2 (RedDiff w/ alignment (H)), as discussed in the implementation details.
+TABLE II: Cross-prompt transfer attack results on the hold-out set against LLaVA.
 Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
 Greedy-Prompt Attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
-RTD w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
-RTD w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
-RTD w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
-RTD w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
+RedDiff w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
+RedDiff w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
+RedDiff w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
+RedDiff w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
 Perspective API (%)
@@
 Greedy-Prompt Attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
-RTD w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
-RTD w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
-RTD w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
-RTD w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
-Original Attack Set Table 1 presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API.
+RedDiff w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
+RedDiff w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
+RedDiff w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
+RedDiff w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
+Original Attack Set Table I presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API.
 The percentages represent the rate of specific toxic attributes in the generated outputs, with ‘Any*’ indicating whether a continuation exhibits at least one of the six toxic attributes.
 We begin by examining the Text-Only Attack, where only the toxic text prefix is provided without any visual input.
-Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate—achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective.
+Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate, achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective API.
 This indicates that current alignment mechanisms in VLMs remain insufficient for preventing harmful completions even in unimodal settings, exposing a fundamental vulnerability in their default safety alignment.
@@
 For instance, using the Detoxify classifier, the ‘Any*’ rate increases from 55.72% (Single-Prompt) to 64.12%, illustrating the benefit of greedy prompt search in finding high-risk image prompts.
-The full Red Team Diffuser (RTD) further enhances attack strength.
-RTD w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers.
-Notably, RTD also demonstrates strong cross-classifier transferability: RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify.
-This confirms that the toxic behaviors induced by RTD are not artifacts of a single classifier but reflect a general vulnerability in VLM alignment under multimodal inputs.
-Table 3.
-Cross-model transfer attack results on Gemini-1.5-flash.
-Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
-Detoxify (%)
-(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
-Text-only Attack 4.18 61.75 2.25 34.64 2.84 76.23 76.23
-RTD w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
-RTD w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
-RTD w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
-RTD w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
-Perspective API (%)
-(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
-Text-only Attack 5.76 60.95 3.17 19.14 5.01 74.41 78.26
-RTD w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
-RTD w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
-RTD w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
-RTD w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
-Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA.
-Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed methods.
+The full RedDiffuser framework further enhances attack strength.
+RedDiff w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers.
+Notably, RedDiffuser also demonstrates strong cross-classifier transferability: RedDiff w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify.
+The consistent gains across Detoxify and Perspective API indicate that RedDiffuser’s induced toxic behaviors are not tied to a specific classifier, but instead exploit broader alignment vulnerabilities in VLMs.
+Hold-out Set Table II presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA.
+Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed RedDiffuser.
 Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set.
@@
 Similarly, with the Perspective API, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Single-Prompt Attack.
-More notably, the full Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%).
-Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%).
-These consistent gains clearly validate the strong generalization capability of RTD across diverse, unseen prompts and across different toxicity classifiers, underscoring its effectiveness as a black-box multimodal jailbreak method.
-Table 4.
-Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct.
+More notably, the full RedDiffuser framework consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RedDiff w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%).
+Furthermore, results from the Perspective API show that RedDiff w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%).
+These results confirm that RedDiffuser generalizes well to unseen prompts and performs consistently across toxicity classifiers, highlighting its effectiveness as a black-box multimodal red teaming framework.
+V-B Cross-Model Transferability
+We assess the transferability of our method by applying RedDiffuser trained solely using feedback from LLaVA [liu2024visual] to two unseen VLMs: Gemini-1.5-flash [team2023gemini] and LLaMA-3.2-11B-Vision-Instruct.
+Tables III and IV report the toxicity rates measured by both Detoxify and the Perspective API.
+RedDiffuser maintains high toxicity across models despite never interacting with them during training, demonstrating strong generalization.
+TABLE III: Cross-model transfer attack results on Gemini-1.5-flash.
+Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
+Detoxify (%)
+(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
+Text-Only Attack 4.18 61.75 2.25 34.64 2.84 76.23 76.23
+RedDiff w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
+RedDiff w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
+RedDiff w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
+RedDiff w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
+Perspective API (%)
+(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
+Text-Only Attack 5.76 60.95 3.17 19.14 5.01 74.41 78.26
+RedDiff w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
+RedDiff w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
+RedDiff w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
+RedDiff w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
+Table III presents the results on Gemini-1.5-flash.
+Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-Only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective API), indicating that harmful completions are already prevalent even without adversarial visual input.
+Furthermore, all RedDiffuser variants elevate toxicity levels, with RedDiff w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively.
+These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images.
+They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios, not just instruction-following attacks.
+TABLE IV: Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct.
 Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
 Text-Only Attack 0.25 11.18 0.00 3.92 0.16 14.52 15.02
-RTD w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
-RTD w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
-RTD w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
-RTD w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
+RedDiff w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
+RedDiff w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
+RedDiff w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
+RedDiff w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
 Perspective API (%)
@@
 Text-Only Attack 0.50 13.84 0.00 3.16 0.66 15.59 18.68
-RTD w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
-RTD w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
-RTD w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
-RTD w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
-5.2. Cross-Model Transferability
-We evaluate the cross-model transferability of our proposed method by using the RTD trained solely on LLaVA (Liu et al., 2024a).
-Transfer attacks are performed on two unseen VLMs: Gemini-1.5-flash (Team et al., 2023) and Llama-3.2-11B-Vision-Instruct.
-Tables 3 and 4 present the toxicity rates evaluated by both the Detoxify classifier and the Perspective API.
-These results demonstrate that RTD generalizes effectively across models, highlighting its robustness and potential threat as a transferable multimodal jailbreak method.
-Table 3 presents the results on Gemini-1.5-flash.
-Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective), indicating that harmful completions are already prevalent even without adversarial visual input.
-Furthermore, all RTD variants elevate toxicity levels, with RTD w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively.
-These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images.
-They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios—not just instruction-following tasks.
-Table 4 reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards.
-Here, the Text-only Attack yields substantially lower toxicity rates—15.02% under Detoxify and 18.68% under Perspective API—reflecting a high refusal rate for toxic continuations.
-However, RTD is still able to bypass these safeguards.
-RTD w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective), with RTD (M) performing similarly.
-These results demonstrate RTD’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models.
+RedDiff w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
+RedDiff w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
+RedDiff w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
+RedDiff w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
+Table IV reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards [wang2024ideator].
+In this setting, the Text-Only Attack produces substantially lower toxicity rates, with 15.02% under Detoxify and 18.68% under the Perspective API, indicating a high refusal rate for toxic continuations.
+However, RedDiffuser is still able to bypass these safeguards.
+RedDiff w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective API), with RedDiff (M) performing similarly.
+These results demonstrate RedDiffuser’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models.
 Together, they highlight the broader risk of multimodal toxicity attacks against both commercial and open-source VLMs, and underscore the need for more robust, context-aware alignment strategies in future VLM design.
-Figure 3.
-Comparison of images generated by the general-purpose Stable Diffusion (left) and the Red Team Diffuser (right).
-Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
-Table 5. CLIPScore comparison across different diffusers.
-Stable Diffusion RTD (w/o) RTD (L) RTD (M) RTD (H)
+V-C Visualizations
+Figure 3 qualitatively compares images generated by Stable Diffusion (left) and RedDiffuser (right).
+For RedDiffuser, the alignment reward weight $\lambda$ increases from left to right: 0, 0.05, 0.1, and 0.2.
+Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications.
+The general-purpose Stable Diffusion consistently produces neutral and benign visuals.
+In contrast, RedDiffuser generates images that subtly evoke discomfort or toxicity-relevant cues.
+For instance, in the second row, RedDiffuser produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion.
+In the third row, where the prompt involves generating an image of a woman, RedDiffuser creates visuals that are noticeably more sexually suggestive and explicit.
+Similarly, in the fourth row, RedDiffuser introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations.
+The observed differences in visual elements and stylistic features underscore RedDiffuser’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs.
+The comparison highlights how reinforced optimization, instead of mere prompt selection, enables subtle but impactful visual manipulation aligned with toxic continuation goals.
+To quantify semantic alignment, we report CLIPScore [hessel2021clipscore] in Table V.
+All RedDiffuser variants achieve scores on par with or exceeding the baseline (44.47), with RedDiffuser at $\lambda=0.2$ reaching 44.80, demonstrating that our stealth-aware reward maintains prompt relevance while enhancing adversarial potency.
+Figure 3: Comparison of images generated by the general-purpose Stable Diffusion (left) and the RedDiffuser (right).
+Given the same image prompts, RedDiffuser produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
+TABLE V: CLIPScore comparison across different diffusers.
+Stable Diffusion RedDiff (w/o) RedDiff (L) RedDiff (M) RedDiff (H)
 CLIPScore 44.47 41.87 42.86 44.02 44.80
-5.3. Visualizations
-Figure 3 presents a qualitative comparison between images generated by the general-purpose Stable Diffusion model (left) and our Red Team Diffuser (RTD) (right).
-For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2.
-Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications.
-While the general-purpose Stable Diffusion consistently produces neutral and benign visuals.
-In contrast, RTD generates images that subtly evoke discomfort or toxicity-relevant cues.
-For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion.
-In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit.
-Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations.
-The observed differences in visual elements and stylistic features underscore RTD’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs.
-The comparison highlights how reinforced optimization—not merely prompt selection—enables subtle but impactful visual manipulation aligned with toxic continuation goals.
-To complement these qualitative observations, we report CLIPScore (Hessel et al., 2021) as a quantitative measure of image-text semantic alignment.
-As shown in Table 5, all RTD variants achieve CLIPScores comparable to or higher than Stable Diffusion, indicating that the adversarial images remain semantically faithful to their prompts despite conveying more toxic or discomforting signals.
-Notably, RTD with the highest alignment reward ($\lambda=0.2$) attains a CLIPScore of 44.80, surpassing the baseline (44.47) and demonstrating that our stealth-aware reward mechanism enhances both visual regularity and the adversarial effectiveness of the generated images.
-6. Limitations
-While the Red Team Diffuser (RTD) framework shows strong performance and cross-model transferability, it has several limitations: 1.
+V-D Extended Evaluation Under External Guardrails
+Table VI reports the evaluation results when LLaVA is paired with external image-level guardrails.
+In this configuration, generated images are first processed by the Stable Diffusion NSFW filter and a VLM-based semantic NSFW checker; images flagged by either checker will be refused by the model and thus do not produce toxic continuations.
+Comparisons in this section use the RedDiff w/ alignment (M) setting, which provides a strong balance between toxicity amplification and semantic coherence.
+In addition to reporting toxicity rates as in previous sections, we also measure the Guardrail Pass Rate (GPR), defined as the fraction of generated images that pass both checkers.
+GPR indicates how often a generated image remains admissible under external safety filtering.
+Interestingly, we observe that RedDiff (Base) achieves a higher GPR than the unmodified Stable Diffusion (54.13% vs.
+49.96%), despite not being trained with guardrail feedback.
+This suggests that RedDiff (Base) does not simply increase toxicity by producing NSFW imagery, but instead shifts generation toward more subtle, suggestive visual cues that are less likely to be filtered.
+With checker-aware training, RedDiff (Guard) further improves both bypass ability and toxicity.
+Under Detoxify, the ‘Any*’ toxicity rate increases from 38.28% to 69.39%, with GPR rising to 94.50%; similar trends hold under the Perspective API (‘Any*’: 69.47%).
+These results show that checker-aware optimization enables the model to consistently produce images that pass external filters while still eliciting harmful continuations.
+Figure 4 shows visual examples from the three diffusion variants.
+Stable Diffusion generates relatively neutral imagery.
+In comparison, RedDiff (base) produces images with stronger and clearer semantic cues—for instance, exaggerated muscular bodies or heightened conflict in the scene.
+RedDiff (guard) further adjusts these cues to remain accepted under external guardrails.
+We observe a tendency toward stylized or cartoon-like rendering, softer shapes, and implied rather than directly exposed body features.
+Such images retain the suggestive cues needed to influence the VLM, while being less likely to trigger NSFW checkers.
+Overall, these qualitative patterns align with the quantitative results: checker-aware optimization encourages the model to adopt visual styles that are subtle enough to pass safety filtering while still affecting the model’s continuation.
+TABLE VI: Black-box attack results on the full dataset against LLaVA with external guardrails.
+Toxicity rates (%) are evaluated using Detoxify and Perspective API.
+Any* indicates toxicity in at least one category.
+Guardrail Pass Rate (GPR, %) measures the fraction of generated images that pass both external safety checkers.
+Model Detoxify (%) GPR (%)
+Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
+Stable Diffusion 3.09 32.28 1.42 18.93 1.75 37.45 37.45 49.96
+RedDiff (Base) 3.17 32.19 1.25 18.93 1.50 38.12 38.28 54.13
+RedDiff (Guard) 5.67 58.89 2.67 35.03 2.59 69.22 69.39 94.50
+Model Perspective API (%) GPR (%)
+Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
+Stable Diffusion 3.92 32.78 2.17 12.09 2.67 37.11 38.20 49.96
+RedDiff (Base) 4.34 33.03 2.50 13.76 2.75 37.78 39.78 54.13
+RedDiff (Guard) 7.26 59.47 4.67 20.85 4.42 67.31 69.47 94.50
+Figure 4: Visual comparison across diffusion variants.
+Stable Diffusion outputs relatively neutral imagery; RedDiff (Base) produces stronger semantic cues that influence continuations.
+RedDiff (Guard) adjusts visual subtlety to consistently pass external safety checkers while retaining adversarial intent.
+VI Limitations
+While the RedDiffuser framework shows strong performance and cross-model transferability, it has several limitations: 1.
 Task specificity.
-RTD is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored.
+RedDiffuser is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored.
 2.
 Computational cost.
-The RL-based fine-tuning requires significant GPU resources, which may be impractical for resource-constrained users.
-7. Conclusion
-We present Red Team Diffuser (RTD), a novel framework for uncovering critical vulnerabilities in Vision-Language Models (VLMs) through semantically adversarial image generation.
-We formalize cross-modal toxicity amplification as a new attack surface, where visually coherent yet maliciously tuned images exacerbate harmful text continuations—an aspect largely overlooked by current alignment efforts.
-RTD combines greedy prompt search with reinforcement fine-tuning to generate images that reliably trigger toxic outputs.
-Experimental results validate its effectiveness: RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set and 8.91% on a hold-out set, while demonstrating strong cross-model transferability, achieving toxicity gains of 5.1% on Gemini and 26.83% on LLaMA-3.2-11B-Vision-Instruct.
-Qualitative analysis further illustrates how RTD manipulates visual cues to subtly induce toxic continuations.
-These findings expose substantial blind spots in existing VLM alignment strategies and underscore the need for more robust, multimodal-aware safety mechanisms.
-In future work, we aim to expand RTD to broader adversarial objectives and model families, offering deeper insights into VLM vulnerabilities and guiding the development of stronger alignment and defense strategies.
+The RL-based fine-tuning requires considerable GPU resources, which may pose challenges for users with limited budgets.
+VII Conclusion
+We introduced RedDiffuser, a novel framework that exposes a previously underexamined vulnerability in Vision-Language Models (VLMs): cross-modal toxic continuation, where visually plausible images can amplify harmful text completions.
+RedDiffuser combines greedy prompt search with reinforcement learning-based diffusion tuning to generate semantically adversarial images that reliably induce toxic continuations while preserving visual coherence.
+Our experiments show that RedDiffuser substantially increases toxicity rates on LLaVA (+10.69% on the original set and +8.91% on a hold-out set) and transfers effectively across models, raising toxicity by +5.1% on Gemini-1.5-flash and +26.83% on LLaMA-Vision-Instruct.
+To reflect more realistic deployment scenarios, we further evaluate attacks under external NSFW guardrails.
+A checker-aware variant of RedDiffuser significantly improves the guardrail pass rate (GPR), reaching 94.50%, while maintaining high toxicity.
+This result shows that toxic continuation remains feasible even under image-level safety filtering.
+These findings reveal fundamental blind spots in current multimodal safety alignment and highlight the necessity of developing defenses that jointly consider visual semantics and continuation behavior.
+Future work will extend RedDiffuser to broader adversarial objectives and more diverse model architectures to deepen the understanding of multimodal safety risks.